### v1

Abstract
Quantum systems have started to emerge as a disruptive technology and enabling platforms – exploiting the principles of quantum mechanics via programmable quantum bits (QuBits) – to achieve quantum supremacy in computing. Academic research, industrial projects (e.g., Amazon Braket, IBM Qiskit), and consortiums like ‘Quantum Flagship’ are striving to develop practically capable and commercially viable quantum computing (QC) systems and technologies. Quantum Computing as a Service (QCaaS) is viewed as a solution attuned to the philosophy of service-orientation that can offer QC resources and platforms, as utility computing, to individuals and organisations who do not own quantum computers. This research investigates a process-centric and architecture-driven approach to offer a software engineering perspective on enabling QCaaS - a.k.a quantum service-orientation. To conduct this research, we employed a two-phase research method comprising (a) a systematic mapping study and (b) an architecture-based development, first to identify the phases of the quantum service development life cycle and subsequently to integrate these phases into a reference architecture that supports QCaaS. The systematic mapping study (SMS) process retrieved a collection of potentially relevant research literature and based on a multi-step selection and qualitative assessment, we selected 41 peer-reviewed studies as published evidence to answer three research questions (RQs). The RQs investigate (i) demographic details in terms of frequency, types, and trends of research, (ii) phases of quantum service development lifecycle to derive a reference architecture for conception, modeling, assembly, and deployment of services, and (iii) emerging trends that can shape future research on software engineering for QCaaS. The results identify a 4-phased development lifecycle along with quantum significant requirements (QSRs), various modeling notations, catalogue of patterns, programming languages, and deployment platforms that can be integrated in a layered reference architecture to engineer QCaaS. Emerging and future research efforts focus on topics such as classifying the QSRs, enabling model-driven or low-code quantum service engineering, and integrating large language models in the quantum service development process. In the broader perspective of quantum software engineering (QSE), the results of this SMS specifically focus on quantum service-orientation by providing empirically grounded findings, highlighting patterns for knowledge reuse, and offering a reference architecture as a blueprint to engineer software for QCaaS.
keywords:
[…]
1 Introduction
Quantum Computers (QCs) aim to disrupt the status-quo of computing – replacing traditional systems and platforms that are driven by digital circuits and modular applications – with hardware and software that operate on the principle of quantum mechanics [1]. QCs that rely on quantum mechanics can exploit quantum circuits via programmable quantum bits (QuBits) to control quantum gates (QuGates) for achieving ‘quantum computational supremacy’ over their classical counterparts, i.e., digital computing systems [2]. QC systems are in a phase of continuous evolution, and despite being in a state of their infancy due to hardware limitations or lack of software ecosystem(s), such systems have started to computationally outperform classical computers in certain application areas such as quantum information processing, bio-inspired computing, and simulation of quantum mechanics [2]. Academic research [1] [3] and industrial initiatives led by technology giants such as IBM, Google, and Microsoft [4] [5] are striving to achieve strategic advantages associated with quantum systems and software technologies in a so-called ‘race to quantum economy’. A recent report presented at the World Economic Forum titled State of Quantum Computing: Building a Quantum Economy highlights that by the year 2022, public and private investments in quantum computing technologies totalled \$35.5 billion [6] and global investments in QCs reached up to \$55 billion in 2024 [7]. Despite the strategic capabilities that quantum supremacy in computing can offer, programming, operationalising, and maintaining a quantum computer represent a complex and radically distinct engineering paradigm compared to the challenges of classical computing and software engineering [1]. QC specific challenges may include quantum domain engineering, circuit modelling, algorithm design, and classic-quantum hybrid system deployment. These challenges arise from a lack of systematic processes, limited tool support, and a scarcity of human expertise in quantum software development. To address these challenges and support QC innovations, state-funded projects and global consortiums are proactively funding initiatives such as the Quantum Flagship [8] and National Quantum Initiative [9] to develop software ecosystems, networking technologies, and human expertise for the alleged quantum leap in computing [10].
Research context and motivation: Service-oriented systems support the utility computing model, enabling individuals and organisations to access multiple computing services on a pay-per-use basis without owning or maintaining them [11]. Since their early adoption, pay-per-use service-driven systems have now grown from data storage, video streaming, and entertainment, to resource-sharing applications that represent a multi-billion dollar industry in service economies [12][13]. The ‘as-a-Service’ (aaS) model has provided the impetus to wide-scale adoption of service systems that can offer a plethora of computing services including but not limited to storage, computation, infrastructure, platform, and software to end-users [12]. The aaS model can enable users and developers who can exploit the QC platforms (e.g., processors, memory, simulators) offered by quantum vendors, such as Amazon, Google, and IBM [14]. Quantum Computing as a Service (QCaaS) as shown in Figure 1 is a recent and quantum-specific genre of aaS model that is built on the philosophy of service-orientation of QC, i.e., pay-per-shot at QC resources instead of owning, programming, and/or maintaining quantum computers [15]. Quantum vendors view pay-per-shot as an opportunistic business model to generate revenue streams from their QC infrastructures, where a shot is a single execution of a quantum task/algorithm on a quantum processing unit (QPU). In simple terms, we define QCaaS as a ‘quantum-specific variation of the aaS model that offers quantum hardware and software resources such as processor, memory, algorithms, and simulations that can be utilised by users (via client nodes) by connecting to a remotely deployed quantum computer (i.e., server)’. The QCaaS model can alleviate the need to own or maintain quantum computers and can help software and service developers to rely on existing knowledge and best practices to develop software services that can be executed on QC platforms [16]. Quantum software services allow developers to encapsulate data and computations in loosely coupled, fine-grained modules to execute tasks on QC platforms [15]. As shown in Figure 1, service developers can leverage reusable knowledge and best practices from classical service engineering—such as microservices, architecture, and patterns—to design quantum services efficiently and effectively. The developed quantum services can enable quantum users, i.e., service requesters, to utilise QC resources offered by quantum vendors, i.e., service providers via pay-per-shot QCaaS model. Without a systematic and structured approach and a lack of expertise in QC software quantum service development risks include inefficient Qubit utilisation, suboptimal quantum-classical task allocation, error-prone circuit design etc., ultimately constraining the success and sustainability of QCaaS [17].
Figure 1: Study Focus: A Generic SE view of QCaaS Model
Software Engineering for QCaaS: Quantum software engineering, as a quantum-specific genre of classical software engineering, leverages established concepts such as architecture models as system blueprints and patterns as reusable knowledge to support software service developers in engineering solutions for QCaaS [18]. By applying SE best practices, developers can, for instance, utilise the quantum-classical split pattern to design hybrid applications, map architectural components to quantum circuits, and enable model-driven quantum code generation [19] [20]. In this research, we have used an SE-focused two-phase research method that relies on a systematic mapping study to investigate existing literature and follow the architecture-based development for an incremental design, implementation, and deployment of quantum software services. Systematic mapping studies (SMS) use an evidence-based software engineering approach to systematically identify, analyse, synthesise, and document research trends on a given topic [21]. The academic community aims to exploit QC platforms to advance state-of-the-art and develop prototypes to support research [14], while the vendors pursue a revenue stream as well as validation and testing of their under-development quantum platforms, offered as a service [15]. Synergising research and development on quantum computing with service-orientation can allow researchers and developers to leverage existing design principles, patterns, architectural styles, and modelling languages of service computing to offer QC as a service [16]. In the context of QSE, this research synthesises the available evidence reflected via published research to provide a software engineering perspective on architecting, implementing, and operationalising QCaaS. We conducted this study to answer three research questions that investigate (i) demographics: trends, types, frequency and progression of research (ii) quantum service lifecycle: that can be integrated in a layered reference architecture for conceptualising, modeling, assembling, and deploying services, and (ii) emerging research: to highlight topics for potentially futuristic research on QCaaS. The results indicate that in quantum service-orientation, QSRs, both functional aspects and quality attributes, guide the four-phase service development lifecycle. Software modeling languages (e.g., UML, circuit diagrams) and patterns (e.g., API gateway, quantum-classic split) can be applied within the proposed layered architecture of QCaaS via quantum-classical hybrid computing. Emerging trends include QSR classification, model-driven development, empirically derived patterns, and defined human roles in QCaaS. The main contributions and implications of this research are to:
  • analyse and synthesise the published research as available evidence on principles and practices of software engineering that can be applied to QCaaS that enables QSE for services computing.
  • identify and document service lifecycle phases into a reference architecture, acting as a service development blueprint, to support classical and quantum-specific service-orientation techniques for QCaaS.
  • highlight prominent trends - identifying existing gaps that reflect the dimensions of future research to address emerging challenges and support the development of next-generation QCaaS solutions.
This research represents one of the pioneering efforts to rely on evidence-based software engineering guidelines to investigate QCaaS and its implications on QSE. The study results can have implications for academic researchers who can rely on mapping of existing research and vision for future work to research and develop QCaaS [1][8]. The results can also guide practitioners who can rely on academic references about patterns (reusability), notations (representation), and implementation (prototyping) to architect QCaaS [14].
Technical context of QC and QSE is presented in Section 2. Research method is discussed in Section 3. Results of the study to answer the RQs are presented from Section 4 - 7. Review of the existing research is presented in Section 8. Implications and validity threats of the study are discussed in Section 9. Conclusions and future work are presented in Section 10.
2 Research Context
This section contextualises the core concepts used in our study and elaborates them based on illustrations in Figure 2. First, we overview the fundamental elements of quantum computing and the role of quantum software engineering to operationalise QC systems in Section 2.1, as shown in Figure 2 a). Second, we explain service-orientation from quantum software engineering perspective in Section 2.2 via Figure 2 b). The concepts and terminologies introduced in this section are used throughout the paper.
Figure 2: An Overview of the Quantum Service-Orientation
2.1 Quantum Computing and Quantum Software Engineering
2.1.1 Quantum Computing Systems
We discuss the quantum computing system from both the hardware and software perspective as shown in Figure 2 a). Figure 2 is adopted from [19] and extended to highlight software architecting in the context of quantum service-orientation. We briefly overview a QC system that comprises of quantum hardware and software elements as shown in Figure 2 a). Fundamental to enabling quantum computation are Quantum Bits (QuBits) that represent the basic unit of quantum information processing by manipulating Quantum Gates (QuGates) [2] [10]. Traditional Binary Digits (Bits) in classical systems (i.e., digital computers) are represented as [1, 0] where 1 represents the computation state as On and 0 represents the state as Off to manipulate binary gates in digital circuits. In comparison, a QuBit represents a two-state quantum computer expressed as $|0\rangle$ and $|1\rangle$. The state of a single QuBit can be expressed as $|0\rangle=\begin{bmatrix}1\\ 0\end{bmatrix}$ and $|1\rangle=\begin{bmatrix}0\\ 1\end{bmatrix}$ and quantum superposition allows a QuBit to attain a liner combination of both states:
$|0\rangle=\left[\begin{array}[]{c}1\\ 0\end{array}\right]~~~~~+~~~~~|1\rangle=\left[\begin{array}[]{c}0\\ 1\end{array}\right]$ (1)
According to Figure 2 a), we distinguish between a Bit and QuBit such that a Bit can take a value as either ‘Off:0’ or ‘On:1’ with 100% probability. In comparison, a QuBit can be in a state of $|0\rangle$ or $|1\rangle$ or in a superposition state with 50% $|0\rangle$ and 50% $|1\rangle$. In addition, two QuBits can be entangled such that entangled QuBits are linked in a way that observing (i.e., measuring) one of the QuBits, can reveal the state of the other QuBit. Extended details about QuBits and QuGates to develop and operate the QC systems are reported in studies like [1] and [14]. To utilise the quantum computing resources, such as quantum processor and memory, there is a need for control software that can program QuBits to manage QuGates of a QC system. To support systems and services for quantum computing, there is a need for systematic processes that incorporates tools, models, patterns, frameworks, and human expertise as engineering practices to develop quantum software [22] [23].
2.1.2 Quantum Software Engineering
QSE has emerged as a quantum specific genre of classical software engineering and it focuses on unifying the principles and practices of quantum mechanics and practices of software engineering - to design, develop, validate, and evolve quantum age software systems, services, and applications [24] [25]. QSE empowers the role of software designers and developers who can exploit processes that support structured development, architectural models for design visualisation, and patterns as best practices to engineer software that can be executed on QCs. Quantum software systems rely on quantum source code compilers that allow quantum algorithm designers and programmers to develop and execute software for quantum computers. For example, a programmer can use a quantum programming language, such as Q# (by Microsoft) or Qiskit (by Google) and use a quantum compiler to enable programable quantum computations [4] [25]. Software systems capable of managing and controlling quantum hardware find applications in areas including, but not limited to, quantum cryptography, bio-inspired computing, and quantum information processing [1]. However, the scarcity of quantum hardware, the shortage of quantum software professionals, and the high costs associated with owning or maintaining quantum computers are critical factors impeding the commercial viability of quantum computing [2] [3]. Vendors who offer QCaaS platforms view quantum service-orientation as an opportunistic business model that offers QC resources to customers as utility computing model [11] [14].
2.2 Service-Orientation for Quantum Computing
2.2.1 Services Computing
It follows the service-oriented architecture (SOA) style that allows service users to discover and utilise a multitude of available software services that encapsulate computing resources and applications offered by service vendors/providers [12]. Fundamental to the concept of service computing is software service that represents a self-contained, reusable, and platform-independent software module that encapsulates specific functionality or business logic, accessible over a network through standardised protocols, typically as part of SOA [26]. More specific types of services can be web services or microservices that enable Software as a Service (SaaS) model in which software applications are hosted and managed by a provider, accessed by users over the network via a web client, and offered on a subscription basis, eliminating the need for local installation and maintenance [15]. Figure 2 b) shows SOA style-based quantum servicing where a QC user (i.e., service requester) can utilise the QC resources offered by quantum vendors (i.e., service provider) by means of quantum services.
2.2.2 Quantum Services Computing
Despite their benefits, QC systems of today lack the capabilities of executing quantum algorithms wrapped with large and diverse amounts of data, inputs, and outputs [2] [10]. As shown in Figure 2 a), large volumes of data in quantum algorithms require more QuBits and complex QuGates that result in deep quantum circuiting and consequently increased error rate, generally referred to as noisy intermediate-scale quantum (NISQ) [4]. To address the issues like NISQ, the classic-quantum split pattern slices the overall quantum software or application into classical modules (pre/post-processing) and quantum modules (quantum computation) that result in hybrid applications[27]. One of the prime examples of classic-quantum split patterns is Shor’s algorithm which involves quantum computations for finding the prime factors of an integer with its application in computer security and cryptography. Quantum computing vendors such as Amazon, IBM, and Google have started to offer their QC systems and infrastructures to be utilised by individuals and organisations by means of quantum services computing [14]. For example, Amazon Braket allows quantum computing services to build, execute, and simulate quantum software based on pay-per-shot model. Ongoing research and development are focused on offering algorithms, hardware, simulations, and mathematical problems as services for quantum platforms [15].
Classical vs. quantum microservices: In service computing model microservices help software designers and developers to structure software applications, via the application of microservice architecture style, as loosely coupled and deployable modules of code that enable computation and data storage [28]. Microservices systems enable enterprises to adopt the service computing model by migrating or modernising their monoliths to a single application that consists of multiple small services [26]. Industry-leading service providers such as Netflix and Amazon represent typical examples of how microservices architecture helps businesses deliver their core business functionality (e.g., video streaming, online shopping) with increased scalability that efficiently serves millions of customers across the globe. Quantum microservicing is a term that refers to microservices that are deployed and executed on quantum computing platforms. For example, the solution [29] utilises Qiskit an open-source quantum software development kit to develop and deliver quantum computation as a collection of microservices that can be executed on IBM Quantum platform. From an implementation perspective, which primarily involves writing and executing source code, both classical and quantum microservice are similar because they implement an algorithm or contain a module of source code that can be executed on QCs [14]. However, from operational and deployment point of view, which focuses on managing the execution of service on quantum platforms, quantum software requirements need to be fulfilled as implemented functionality and desired quality of the microservice.
We conclude that, in order to harness QC as utility computing, there is a need to tailor existing principles and methods of service-orientation or develop new architectures, frameworks, and empirically grounded processes to synergise QC and SOA in the context QCaaS. QSE can provide the conceptual foundations, architectural blueprints, and reusable patterns necessary to navigate the complexities of QCaaS. By embracing SE principles—such as the quantum-classical split, architectural mapping to quantum circuits, and model-driven code generation—developers can orchestrate hybrid applications that synergise classical and quantum computation. Quantum service-orientation when viewedăfrom a utility computing perspective has the potential to bridge the quantum divide — a pressing issue highlighted at the World Economic Forum 2023 — between states and entities that possess quantum computing systems, technologies, and infrastructure, and those that do not [30]. The QCaaS also provides foundations for building a quantum computing cloud (QC cloud or simply quantum cloud), led by technology giants like IBM to offer computation resources to the end-users based on cloud computing [31].
L Quantum software engineering as a discipline refers to the systematic application of software engineering principles to quantum computations by adapting engineering practices to hybrid classical–quantum environments and to the constraints of emerging quantum hardware. For example, the Qiskit SDK ecosystem demonstrates QSE in practice, as it provides structured frameworks for designing quantum circuits, integrating them with classical workflows, and testing their execution across simulators and real quantum processors. Quantum service computing refers to the application of service-oriented principles to the design, composition, and execution of quantum computing functionalities, where quantum resources and capabilities are encapsulated as interoperable services. For example, Amazon Braket provides qunatum service computing by exposing quantum resources (e.g., circuit simulators and quantum hardware backends via multiple vendors) as services, enabling developers to compose QC tasks. Quantum software service refers to the abstraction and delivery of quantum software capabilities, such as algorithms, compilers, or simulators etc., as modular services that can be executed on quantum platforms (simulators or physical processors) through service-oriented architectures. For example, in the IBM Quantum Experience, algorithms such as Grover’s search are exposed as cloud-based services accessible via APIs like Qiskit, allowing developers to seamlessly invoke and run them on both simulators and actual quantum hardware without managing QC complexities.
3 Research Method
Figure 3: An Overview of the Research Method
To conduct this study, we used a two-phase research method including (a) systematic mapping and (b) architecture-based development, detailed below and illustrated in Figure 3.
  • Systematic Mapping Study (SMS): First, we followed the guidelines and steps of a systematic mapping study that relies on evidence-based software engineering approach to identify, analyse, classify, map, and document available evidence on the topic under investigation [32]. According to the guidelines in [21], a systematic mapping study offers ‘a means of evaluating and interpreting all available research (as existing evidence) relevant to a particular question, topic area, or phenomenon of interest’. To ensure that our research method adheres to the recommendations and guidelines for conducting the mapping study [21], we followed a multi-step process that is illustrated in Figure 3 and detailed in the remainder of this section. It is worth mentioning here that in recent years, there is a growing interest in conducting multivocal literature reviews (MLRs), which incorporate grey literature (often non-published, non-peer reviewed material), especially in fast-evolving fields like quantum computing and quantum software engineering [33]. In this study, we have chosen the SMS approach, adhering to the guidelines of conducting systematic reviews [34], to focus solely on peer-reviewed published research as the primary source of evidence on quantum service computing. Although non-peer-reviewed studies and gray literature such as [16] [35] are considered in this paper to provide additional context for the review of related research (Section 8) that complements the findings of primary studies (Appendix A). Non peer-reviewed and non-published studies are treated as supplementary sources and are not included in the list of primary studies for this research.
  • Architecture-based Development (ABD): As part of SE process, ABD is also referred to as architecture-centric software development [36] where architecturally significant requirements [37] are modelled as system blue-print to guide the subsequent phases of ABD, such as the conceptualisation, implementation, and evaluation of software to be engineered. ABD is generally classified into three main activities, namely architectural analysis, architectural synthesis, and architectural evaluation [38]. The first activity, architectural analysis, is driven by a scenario or architecture story as a description of the problem or potential use-cases for the architecture. The next activity architectural synthesis consolidates the details in the story to create an architecture model or representation that can act as a point of reference, visualising the structural (de-) composition of the software. Finally, the last activity architectural evaluation evaluates the synthesised architecture against ASRs based on scenarios from the architectural story.
In the following, first, the steps of the SMS process are detailed from Section 3.1 to Section 3.5 that follow ABD, which is discussed in Section 3.5.
3.1 Step 1 - Specifying the Research Questions
As the initial step of the research method, we outlined three research questions (RQs) whose answers represent the results of this research. While formulating these RQs, as per the guidelines of systematic mapping studies [21] and systematic reviews [39], we focused on three key aspects. These include i) analysing the demography of published research, ii) synthesising the collective evidence on quantum service engineering, i.e., service lifecycle activities, and iii) highlighting prominent trends of emerging challenges/solutions. Each of the RQs is elaborated by discussing the core objectives associated with each question.
RQ-1: What are the demographic details of published research on software engineering for quantum computing as a service?
Objectives - The demographic details can include diverse information regarding the frequency, types, core contributions of available evidence, i.e., existing published research. Such details help researchers to identify emerging and potentially futuristic trends of publications on quantum service computing. We specifically focus on four key aspects of demographic details that include the frequency of publications, types of publications, types of published research and application domains (e.g., use-cases, scenarios) to which research-based solutions have been applied in the context of QCaaS.
RQ-2: What software engineering solutions are reported in the literature to enable or enhance quantum computing as a service?
Objectives - Software engineering solution refers to SE artifacts that include but are not limited to process, patterns, architectures, frameworks, tools, etc. that support a systematic and incremental approach to architect, implement and/or reengineer software services. A review of SE solutions indicates the extent to which classical or quantum specific software engineering theory and practice can be applied to engineer QCaaS [31]. We have focused on identifying the SE solutions and aligning them with four activities of service lifecycle including functional and quality aspects (service conception), notations and patterns (service modeling), implementation and platforms (service assembly), and execution platforms (service deployment) to support QSE in the context of service development for QCaaS. SE solutions and service lifecycle activities provide the foundation to derive a reference architecture for QCaaS.
RQ-3: What are the emerging trends of research on software engineering for quantum computing as a service?
Objectives - Emerging trends indicate open challenges or potentially futuristic research that can shape the agenda, provide a road map for progressing research and development, and outline new hypotheses to be tested for QCaaS. We have identified the trends that can help pinpoint the prevalent challenges and their solutions to help researchers and practitioners in developing solutions for emerging and next-generation of QCaaS.
Table 1: Criteria for Screening and Qualitative Assessment of Selected Studies
Study Selection Step I - Screening of Identified Studies
S1 - The study does not discuss any solution or proposal for quantum computing as a service
S2 - The study is not reported in English
S3 - The study is a duplicate study. Duplicate studies are studies with overlapping contents, e.g., a conference paper extended as a journal article.
S4 - The study is a secondary study/survey paper
Exclude the study if the answer to any of the criteria in Step I (S1 - S4) results in Yes, otherwise,
Include the study for quality assessment in Step II
Study Selection Step II - Quality Assessment of the Identified Studies
Q1 - Study objectives and Contributions are clear? [Yes = 1, Partially = 0.5, No = 0]
Q2 - Research method to conduct the study is reported [Yes = 1, Partially = 0.5, No = 0]
Q3 - Design and/or implementation details of solution are provided [Yes = 1, Partially = 0.5, No = 0]
Q4 - Details for Experiments/Evaluation/Demonstration of Solution are provided [Yes = 1, Partially = 0.5, No = 0]
Q5 - Study limitations and needs for future research are discussed [Yes = 1, Partially = 0.5, No = 0]
Exclude the study that has a quality assessment score (Q1 – Q5) less than 2.0, otherwise
Include the study for review and data extraction in Appendix A - Table 4
3.2 Step 2 - Search String and Electronic Data Sources
The initial identification of key terms or keywords for literature was carried out by the first three authors through an analysis of the RQs. The goal of identifying the key research terms was to formulate a search string that can be customised as per specific Electronic Data Sources (EDS) for an automated search, as per the guidelines in [40]. Ultimately, the authors reached a consensus on the search string in Listing 1. The key terms were combined using the boolean operators “OR” and “AND” to formulate the search string. This final search string was developed following a pilot search of relevant literature on IEEE Xplore and Google Scholar. The purpose of the pilot search was to uncover existing study titles and various synonyms for software engineering and quantum software engineering in the context of software services. For example, we discovered that using the term ‘framework’ instead of the term ‘process’ returned a significant number of irrelevant studies focused on quantum computing frameworks that were related to QC in general with no relevance to QSE or quantum service computing. Moreover, inclusion of key-terms such as ‘qubit’ retrieved hardware-focused studies, not software and/or service models, thus adding irrelevant studies (noise). We also dropped the key terms ‘quantum–classical’ and ’quantum classic’ as both overlapped with the main term ’quantum’ that alone sufficiently retrieves most relevant studies and avoids an extensive list of irrelevant studies to be scanned.The primary goal of the final search string was to identify the most relevant literature while minimizing the inclusion of irrelevant studies, thereby leading to a manual review of titles, keywords, and abstracts for study selection.
Listing 1: Search String to Identify the Primary Studies
"Quantum" AND
("Service" OR "Cloud" OR "Utility" OR "Microservice" OR "MSA") AND
("Compute" OR "Process" OR "Platform" OR "Infrastructure" OR "Environment")
When conducting systematic reviews and systematic mapping studies, EDS facilitates automated searches using predefined and often customised search strings to identify relevant literature about the topic under investigation [41]. Following recommendations for implementing a systematic search process and choosing the most appropriate data sources, we selected five EDS for automated search [42]. These include IEEE Xplore, ACM Digital Library, Springer Link, ScienceDirect, and Wiley Online Library, which are widely recognised as key sources for literature on computing, particularly software engineering and service computing research. While this list is not exhaustive and does not claim to cover all existing literature, prior SLR or SMS-based studies have identified these five electronic sources as sufficiently comprehensive and appropriate for indexing relevant research studies or evidence.
The search process included execution of the search string in Listing 1 that was customised in the context of each EDS in Figure 3. As a specific example, we utilised the advanced search feature which is available on IEEE Xplore including the ‘Search Term’ option to execute the search string and find published studies within ‘Full Text & MetaData’ To refine our results and to remove potentially irrelevant studies, we adjusted the search parameter, switching from ‘in Full Text & MetaData‘ to ‘in Abstract.’ The search conducted between December 10 - January 3, 2025 included all published studies until December 2024, which resulted in a retrieval of titles of 287 research articles representing potential primary studies. As part of this activity, the search results were subsequently filtered by the first two authors based on titles, keywords, and abstracts by following inclusion and exclusion criteria, i.e., study screening in Table 1 that resulted in limiting the relevant studies to 105 studies. Finally, a full-text analysis was conducted, leading to the selection of 55 primary studies for data extraction. Third, fourth and fifth authors were invited to review the obtained search results and the list of selected articles.
Following the automated search via EDS, we employed the snowballing technique to manually examine the reference lists of the 55 selected primary studies, aiming to uncover any additional studies that may have been missed during the initial string-based review process [43]. The snowballing technique led us to identify two more studies that were screened for inclusion or exclusion. The first and second authors primarily conducted the snowballing technique, while the third and fourth authors were invited to collaboratively verify their findings to avoid potential bias.
3.3 Step 3 - Study Selection - Primary Studies for the SMS
To select the primary studies for SMS, we followed the recommendations by Kitchenham et al. [39] for including or excluding identified studies. Table 1 presents a two-step criteria-driven process for literature inclusion and exclusion via screening and quality assessment. The first step, referred to as screening, enabled us to filter out any irrelevant, redundant, or non-English studies. The second step, namely quality assessment, enabled us to do a qualitative assessment of each included study, eliminating those that did not meet the formulated quality criteria. The established criteria effectively filtered the search results produced by the search string.
3.3.1 Searching the Primary Studies
Based on automated search and snowballing, a total of 41 studies were ultimately selected for review and analysis to answer the RQs, as in Figure 3. The complete list of these primary studies is detailed in Appendix A - Table LABEL:tab:study-list.
3.3.2 Performing Quality Assessment
The selected studies were evaluated using quality assessment criteria from Table 1 to minimise research bias and assess the significance and completeness of each study. These criteria includes five specific assessment questions (Q1 – Q5), against which each primary study was evaluated. A score of 1 was assigned if a study explicitly addressed the specific QA, 0.5 if the QA was partially addressed, and 0 if there was no evidence of adressing the QA. The total quality score for each study was calculated by summing up the scores assigned for all the QAs. The first author conducted the quality assessment, and the second and third authors independently verified the results for accuracy and consistency. Studies with a cumulative QAs score of 1.5 or higher were included in the final list.
3.4 Step 4 - Extracting Data from Primary Studies
A set of data extraction items were defined to extract specific data related to the RQs. These items represent specific data points extracted from each primary study, aligned with the RQs. The template for data extraction is presented in Appendix A - Table 4. The first author conducted a pilot extraction on 10 studies to evaluate the appropriateness of the data items, which were revised based on feedback from the second and third authors. The first three authors collaboratively performed the formal data extraction. The extracted data was synthesised before documenting the results of the SMS. The extracted data is available via [44].
3.5 Step 5 - Documenting the Mapping Study
The last step involved documenting the results of SMS, representing answers to individual RQs in dedicated sections. Specifically, the results of the SMS documented as demographic details of published research are in Section 4 (answering RQ-1), software engineering solutions are discussed in Section 5 - 6 (answering RQ-2), and emerging trends are highlighted in Section 7 (answering RQ-3).
Supplementary material for the study is provided as part of the extended details of the documentation. Specifically, the extracted data from primary studies - organized according to the data extraction form presented in Appendix A forms the basis for answering the RQs and documenting the SMS, and is made available via [44]. Moreover, a dedicated GitHub repository and a short video demonstration showcasing the architecture-based development of quantum software are accessible at [45] and [46], respectively.
4 Demography of Published Research (RQ-1)
This section overviews the demographic details of published research to answer RQ-1 based on illustrations in Figure 4. Specifically, RQ-1 presents demographic details about the years and types of publications (Section 4.1), type of research (Section 4.2), and application of research solutions to specific domains (Section 4.3) that complement the answers to other RQs. Demographic details can reveal trends such as most recent research, e.g., journal articles [S3][S33] [^1] The notation [SN], where S = study and N = an integer between 1 - 41 such as S1 (first study in Appendix A - Table LABEL:tab:study-list) indicates a unique reference in the list of primary studies selected for the SMS. This notation also helps to distinguish between the selected primary studies and references in the bibliography section of this paper. published in 2024 propose an innovative solution or evaluate existing research to engineer quantum software for enabling qunatum cloud computing and computational optimisation, further elaborated below. Throughout this paper, while answering the RQs, a summary of the core findings and guidelines representing key takeaway message(s) is provided which is denoted with symbol L.
Figure 4: Demographic Details of Published Research on QCaaS
4.1 Frequency and Types of Publications
Analysing the frequency and types can highlight the relative growth and trends of publications that indicate a) number of peer-reviewed published studies (i.e., frequency: publications per year) and b) prominent fora of publications (i.e., publication types: journal article, conference proceedings etc.) on topic under investigation. The illustrations in Figure 4 a) and Figure 4 b) visualise the core findings about frequency and types of publications discussed below. Specifically, Figure 4 a) shows four years including 2021, 2022, 2023, 2024, presented along the horizontal-axis, and number of publications in a given year, represented on vertical-axis. It is vital to mention that Figure 4 indicates no published study prior to the year 2021 and the reason for this can be two-fold including (i) our literature search (Section 3, Figure 3) did not identify any relevant studies before 2021, or (ii) any identified study before 2021 did not pass the quality evaluation criteria (Table 1) for inclusion in the list, i.e., Appendix A. The studies [S5] [S2] published as workshop paper and conference proceedings in 2021 represent one of the earliest published research on experimenting with microservices that can be executed on quantum computing platforms. The most recent research, for example as reported in [S10][S18] published in 2024 as journal articles focuses on architecting software services for QC platforms. The bar graph reveals a significant growth of published research in the year 2023 and 2024 (i.e., 29/41 studies, reflecting 71% approx.) of the selected studies highlighting progressing of research from experimental solutions [S5, S9] to identifying patterns [S16] and creating reference architecture [S10] that support service engineering for QC software.
Figure 4 b) indicates five types of publications, also referred to as publication fora classified as Journal articles (11/41 studies, i.e., 27%), Conference proceedings (19, 46%), Symposium papers (4, 10%), Workshop Papers (5, 12%), and Book Chapters (2, 5%). Conference proceedings represent the most prominent fora of publications with research focused on publications on quantum function as a service [S3], exploring quantum service-oriantation and cloud computing [S20], quantum microservicing for smart systems [S25], and quantum-classic hybrid computing [S32]. A 2021 report by Scopus also highlights emerging research trends in quantum computing, based on Scopus-indexed documents, with a focus on the demographic details of published research in general area of quantum computing systems, software, and services. The report offers a comprehensive overview, analysing the publication focus, leading institutions, and key contributors in the field of quantum computing [47].
L Publication frequency There is a significant growth in published research in last 2 years: 2023-2024 (n = 29, 71%) that propose emerging solutions such as quantum-classic hybrid software, quantum cloud and edge computing as well as leveraging classical service engineering techniques such as quantum SOA, quantum web servicing that can be customised or reused for quantum software systems. Publication types Conference proceedings and journal articles are most prominent fora of publications disseminating research via well-established and top-ranked software and service research venues, e.g., JSS, FSE, ICSOC, IEEE Cloud. Some recently organised workshops, e.g., QP4SE conference series, e.g., IEEE QSW and edited book, Quantum Software Engineering represent venues dedicated to publishing theory and practice of quantum software and services.
4.2 Types of Research
The type of research, also commonly referred to as research type facet classifies the published studies based on their specific contribution, such as an innovative model, architecture, framework, or prototype as a newly proposed solution [S1, S3, S10] or validation of an existing solution with a new dataset or case studies [S15]. The classification for the type of research is based on six research type facets namely experience report, evaluation research, opinion papers, philosophical papers, solution proposal, and validation based on classification criteria in [21]. Our analysis did not find any experience report that documents personal experiences or perspectives of researchers and practitioners on a specific problem and/or project, shown in Figure 4 c). Solution proposal is the most prominent type of published research (26 studies, 63%) that document solutions such as processes [S26, S33], patterns [S16, S18], architectures [S10], frameworks [S3, S1], and prototype [S29] etc. for quantum software services. In contrast, we only identified one philosophical paper [S8] that argues about hybrid nature of quantum services that split computation as classical and quantum components. The paper also sketches a toolchain to synergise classical and quantum computations into a hybrid computing model. Evaluation research (4 studies), opinion papers (4 studies), and validation research (6 studies) are reported in a total of 14 studies, representing a total of 34% approx. Specifically, evaluation research aims to evaluate an existing solution based on experimentation, such as a case study on quantum service-orientation in smart cities [S11] and evaluating system performance via simulation of distributed quantum computing [S13]. Opinion papers discuss authors’ negative or positive opinions, i.e., pros and cons regarding a specific technique such as orchestration [S18], deployment [S15], and cloud-based [S19] quantum servicing. Finally, the validation research aims to validate the stated hypothesis, objectives, or proposed features of the solution, which have not yet been deployed in a real-world environment. These types of studies focused on validating the integration of quantum-classic microservices [S5], support of quantum algorithm as a service [S6] and enabling the quantum SOA [S7].
L Type of research Analysing the types of published research highlights that solution proposals represent the most prominent type of published research with a total 26 out 41. QCaaS as a recent and continuously evolving phenomena, i.e., subdomain of QSE relies on theory and practices of service-orientation to propose solution for quantum service-orientation. The role of validation and evaluation type research (total of 8 studies) can be vital to empirically assess existing solutions and methods that can be applied to emerging challenges of QCaaS.
4.3 Application Domains
Application domain refers to the usage scenarios, usecases, or system domains to which the proposed research solutions or findings such as quantum service-orientation and cloud computing [S22, S25] or qunatum API gateway can be applied [S1]. For example, the research in [S11, S23] proposes a solution of Quantum as a Service (QaaS) to secure a smart city systems. This refers to quantum service-orientation (i.e., proposed research) applied to smart systems (i.e., application domain). As illustrated in Figure 4, our analysis classified the existing research into seven research domains including a) cyber security (3/41 studies, 7% approx.) b) computation optimisation (14, 34%) c) experimental development (5, 12%) d) quantum service orientation and cloud computing (10, 25%) e) smart systems (2, 5%), f) simulation (3, 7%), and (g) others (4, 10%), each exemplified below. Computational optimisation and qunatum service and cloud computing represent the most prominent domains with a total of 24 studies, 59% of total studies. For example, the studies [S14, S28] focus on optimising the data search problem with quantum search algorithms. Similarly, the studies [S7] focuses on service orientation of quantum resources and enabling the quantum cloud computing [S19, S22]. We found a total of 4 studies that can be classified under others including traffic flow [13] and 5G networking [S40].
L Application domain Research-based solutions on quantum services computing has diverse application domains ranging from optimisation, simulations, smart systems, and quantum clouds. Quantum cloud computing as an emerging domain synergises the cloud computation model with quantum software services to offer QCaaS that can be applied to diverse domains including but not limited to smart systems and cybersecurity solutions.
5 Solutions to Engineer Software for QCaaS (RQ-2)
We now answer RQ-2 that aims to investigate the existing SE methods, as published research, to enable or enhance quantum computing as a service. For a fine-grained presentation and systematic interpretation of RQ-2 results, we present the data from reviewed studies based on the following four points and summary of results in Table 2 and visualisation of core finding in Figure 5. The results can be interpreted using:
  1. 1.
    Service lifecycle: Four activities namely conception, modelling, assembly, and deployment are collectively referred to as foundational activities for IBM SOA lifecycle [26]. These four generic activities are used to contextualise and organise seven steps to engineer software services for quantum computing. Specifically, an activity guides about what needs to be done? whereas, the step illustrates how it is to be done? For example in study [S1], the first activity namely conception provides foundations for system analysts to conceive the idea of developing a service to execute mathematical computations. Functional aspects and quality attributes as two specific steps of conception can guide service developers to consider the required functionality (e.g., quantum service delivery) and quality (e.g., dynamic adaptation) that needs to ensured in the implemented service. Each of the seven steps to engineer QCaaS are organised under four activities of SOA lifecycle and presented in Table 2.
  2. 2.
    Extracted data from studies: The data extracted from primary studies recorded in represents the published evidence that is synthesised and summarised in Table 2. Table 2 acts as a structured catalogue, detailing each reviewed study that represents published evidence (Study ID) and seven steps of QCaaS under four service lifecycle activities. For example, Table 2 can be looked up to quickly identify that study referred to as [S1] focuses on (i) conception as optimal delivery of quantum service (functional aspect) via a runtime dynamic adaptation and selection of a specific service (quality attribute) from a pool of available services. To support (ii) modelling of this service UML deployment diagram is used as a notation and API Gateway pattern is applied to manage client request and runtime selection of most suitable service. To (iii) assemble this service, a usecase on process automation and service optimisation is selected and Flask as a Python-based framework is used to implement the service. Finally, for (iv) deployment of the service Amazon Braket is used as a service delivery platform.
    Table 2: Summary of the Phases and Activities of Quantum Service Lifecycle (identified from the reviewed literature)
    Published Evidence Conception Modeling Assembly Deployment
    Study ID Functional Aspects Quality Attributes Modeling Notation Design Pattern Service Usecase Programming Language Service Platform
    S1 Quantum Service Delivery Dynamic Adaptation UML API Gateway Pattern Process Automation and Optimisation Python,Flask Amazon Braket
    S2 Quantum-Classic Hybrid Computing Service Integration and Intreoperability Process Models Miscelleneous Process Automation and Optimisation N/A N/A
    S3 Miscelleneous Dynamic Adaptation Unified Modeling Language Quantum-Classic Split Experimental Computation Microsoft technology, Q# IBM Quantum
    S4 Quantum Experimental Services Miscelleneous Unified Modeling Language No pattern Experimenatal Computation Python Amazon Braket
    S5 Quantum Experimental Services N/A No explicit notation Service Composition Process Automation and Optimisation Python,Flask Amazon Braket
    S6 Quantum Experimental Services Service Integration and Intreoperability UML API Gateway Pattern Algorithmic Exectuion Python Rigetti
    S7 Quantum Experimental Services Service Integration and Intreoperability Graph Models API Gateway Pattern Process Automation and Optimisation N/A Amazon Braket
    S8 Miscelleneous Service Integration and Intreoperability Graph Models Miscelleneous Quantum Simulation N/A N/A
    S9 Quantum Experimental Services Service Performance No explicit notation Service Composition N/A Python Amazon Braket
    S10 Quantum Service Delivery Usability UML Quantum-Classic Split Cyber Security C# Microsoft Azure Quantum
    S11 Quantum Service Delivery Security Miscelleneous Service Composition Smart City Qiskit, D-Wave libraries IBM Quantum
    S12 Quantum-Classic Hybrid Computing Dynamic Adaptation Miscelleneous No pattern Process Automation and Optimisation OpenAPI N/A
    S13 Quantum Service Delivery Service Performance UML Miscelleneous Process Automation and Optimisation N/A N/A
    S14 Quantum Service Delivery Miscelleneous UML No pattern Cyber Security N/A IBM Quantum
    S15 Quantum Service Delivery Usability UML API Gateway Pattern Process Automation and Optimisation Python, Flask Google Quantum AI
    S16 Quantum Cloud and Service Orchestration Computational Efficiency Process Models Quantum-Classic Split Experimental Computation Python, Zato Google Quantum AI
    S17 Continuous Deployment Computational Efficiency Process Models No pattern Quantum search Python,Flask Amazon Braket
    S18 Quantum Cloud and Service Orchestration Computational Efficiency Miscelleneous API Gateway Pattern Quantum search Python,Flask IBM Quantum
    S19 Quantum Service Delivery Computational Efficiency Miscelleneous Miscelleneous Quantum Simulation Python,Flask IBM Quantum
    S20 Quantum Service Delivery Reliability No explicit notation Service Facade Quantum Simulation Python IBM Quantum
    S21 Continuous Deployment Security Workflow model API Gateway Pattern Algorithmic Exectuion Qiskit N/A
    S22 Quantum Service Delivery Computational Efficiency No explicit notation No pattern Process Automation and Optimisation N/A IBM Quantum
    S23 Quantum Service Delivery Security Workflow model API Gateway Pattern Process Automation and Optimisation Python D-Wave
    S24 Quantum-Classic Hybrid Computing Computational Efficiency Miscelleneous No pattern Experimental Computation Python, and open sources libraries Amazon Braket
    S25 Quantum Service Delivery Computational Efficiency Miscelleneous Service Composition Experimental Computation Qiskit Amazon Braket
    S26 Quantum Service Delivery Computational Efficiency Miscelleneous Service Facade Experimental Computation Python, Flask, D-Wave libraries Amazon Braket
    S27 Quantum Service Delivery Computational Efficiency Miscelleneous No pattern Experimental Computation N/A Amazon Braket
    S28 Quantum Classic Hybrid Computing Computational Efficiency Miscelleneous N/A N/A N/A N/A
    S29 Quantum service delivery Computational Efficiency Workflow model No pattern Experimental Computation Python Amazon Braket
    S30 Quantum Cloud and Service Orchestration Usability Workflow model No pattern N/A N/A N/A
    S31 Quantum Service Delivery Usability Miscelleneous Service Facade N/A Python, D-Wave’s Ocean) Amazon Braket
    S32 Quantum Service Delivery Useablity Miscelleneous No pattern Process Automation and Optimisation Python IBM Quantum
    S33 Quantum Service Delivery Reliability Miscelleneous Service Composition Process Automation and Optimisation N/A N/A
    S34 Quantum Service Delivery Computational Efficiency Workflow model No pattern Process Automation and Optimisation Python, Flask, RESTful API Microsoft Azure Quantum
    S35 Quantum Service Delivery Computational Efficiency Miscelleneous No pattern Quantum Simulation Python N/A
    S36 Quantum Cloud and Service Orchestration Computational Efficiency No explicit notation No pattern Process Automation and Optimisation Qiskit IBM Quantum
    S37 Quantum Service Delivery Computational Efficiency No explicit notation No pattern Process Automation and Optimisation N/A N/A
    S38 Quantum Cloud and Service Orchestration Computational Efficiency Miscelleneous No pattern Process Automation and Optimisation N/A N/A
    S39 Quantum Service Delivery Computational Efficiency Miscelleneous No pattern Process Automation and Optimisation N/A N/A
    S40 Miscelleneous Security cloud Miscelleneous No pattern Security N/A N/A
    S41 Continuous Deployment Computational Efficiency No explicit notation Service Composition Process Automation and Optimisation N/A D-Wave Leap
  3. 3.
    Summary of core findings: Figure 5 accumulates a visual representation of the data from Table 2 via Figure 5 a) to Figure 5 g) and highlights key findings and prominent insights from each of the seven steps for developing QCaaS. For example, Figure 5 d) shows that patterns used to model QCaaS can be generally classified into five types, namely API Gateway, Service Composition, Quantum-Classic Split, Service Façade, Miscellaneous patterns such as enterprise service bus and layered architecture. A total of 17 out of 41 studies (i.e., 42% of reviewed studies) do not explicitly specify any pattern. Service Façade is applied to a total of 8 studies to support service integration and wrapping of classical services to the interfaces of quantum services. Figure 6 complements the discussion on quantum service patterns by providing a thumbnail view, as a visual reference, for the identified patterns.
  4. 4.
    Illustrative example: We also exemplify Shor’s algorithms as an example to architect quantum service-orientation. The example provides a scenario to and a step-by-step illustration and elaboration of the four phases and seven activities of service lifecycle to engineer QCaaS, presented in Figure 8.
5.1 Activity I - Service Conception
The service conception activity focuses on conceptualising the required functionality and desired quality of the service(s) to be implemented. From service engineering perspective, functional aspect and quality attributes that have a profound impact on the design, implementation, and delivery etc. of software services are collectively referred to as architecturally significant requirements (ASRs) [48]. In the context of quantum service-orientation, ASRs also referred to as quantum significant requirements (QSR) that are crucial for the design, development, execution, or integration of QC capabilities within software-intensive systems, applications, and services. QSRs ensure that software can effectively leverage functionality and quality, such as quantum algorithmic implementation, interaction with quantum hardware, or maintain compatibility with hybrid quantum-classical environments. For example, while architecting quantum services [14] [15], QSRs may require a split of a quantum algorithm in terms of classical and quantum computations for executing data inputs, state preparation, and simulations on a combination of classical and quantum processors that represents required functionality. Moreover, such an execution requires efficient utilisation of QuBits and quantum error minimisation as the desired quality.
5.1.1 Functional Aspects
Based on our review, we identified and classified functional aspects into five (5) specific types while quality attributes are classified into seven (7) specific types as highlighted in Figure 5 a). These five types include Continuous Deployment (identified in a total of 3 out of 41 (3/41) reviewed studies that represent 7% approx. of reviewed literature), Quantum Experimental Services (5/41, 12.5%), Quantum Cloud and Service Orchestration (5/41, 12.5%), Quantum-Classic Hybrid Computing (4/41, 10%), Quantum Service Delivery (2/41, 51%), and Miscelleneous (3/41, 7%). In the context of quantum service patterns, some stuides such as [S10][S16] refer to the qnatum-classic hybrid comoputing as qunatum-classic split pattern. For the sake of clarity, the quantum-classic hybrid computing is a general reference to the computing model, whereas quantum-classic split is a specific pattern to implement the hybrid computing model. In the following, we briefly introduce each specific type of functional aspects and quality attributes and exemplify one of them as a representative example for elaboration. For example, considering ‘quantum cloud and service orchestration’ type, the studies [S32][S38] focus on orchestrating software services that can manipulate QC hardware that is deployed on a remote server and made available via the cloud computing model. Specifically, [S38] allows QC vendors to offer quantum resources (e.g., QuBits) that can be utilised via on-demand plans and pay-per-use model, as illustrated in Figure 2. As part of service conception, we summarise the functional aspects as:
  • Continuous deployment enabled via quantum micro servicing is identified in a total of 3 out of 41 studies (representing 7% of all reviewed studies). By following continuous deployment, the functionality of a service is developed, tested, and deployed for a seamless delivery of software services that can be executed on QC platforms.
  • Quantum experimental services support proof-of-the-concepts or experimental delivery of software, including but not limited to enabling number crunching, experimental optimisation, or algorithmic execution on a QC platform reported in a total of 3 studies. Experimental service development offers only a limited functionality, merely capable of checking the status of quantum circuits or generation of random numbers using quantum hardware, referred to as number crunching with QCs.
  • Quantum cloud and service orchestration enables end-users to invoke software services using cloud computing model identified in a total of 5 studies. Quantum cloud computing is gaining significant attention from leading QC vendors such as IBM cloud (IBM quantum experience) [49] and Google Cloud (Google quantum AI) [50] with a push for experimentation with quantum algorithms via cloud-based quantum processors.
  • Quantum-Classic hybrid computing focuses on a hybrid computation model that splits a software into two parts namely classical software and quantum software, presented in a total of 4 studies. The hybrid computing model ensures that quantum processing is utilised only where it offers an advantage (so-called quantum supremacy [2]), while classical computer can handle the routine (classic computation) operations efficiently. In the context of hybrid computing model, we further elaborate on the quantum-classic split pattern based on illustration in Figure 6.
  • Quantum service delivery focuses how software is provided, maintained, and accessed by users, whether through SaaS, on-premises, or managed services for QCs, as identified in a total of 21 studies. Conceptually similar to the concepts of continuous deployment and services orchestration, quantum service delivery is different because it focuses on how users access and interact with software services rather than the technical mechanisms of deployment. In contrast, Continuous Deployment (CD) is a DevOps practice that can enable automation of the release of software updates without manual intervention, ensuring frequent improvements to service delivery.
  • Miscellaneous represents a multitude of functional aspects that cannot be classified under a specific type and are identified in a total of 3 studies. These include, for example, the studies [S3] demonstrating the efficiency of data search via QC, [S8] discussing QC applications for bio-inspired computing, and [S40] discussing quantum-safe security. For example, the study [S8] provides a reference model for QC toolchain and its application to address challenges of bio-inspired computing in the context of humanities and natural sciences.
5.1.2 Quality Attributes
Quality attributes for software services define the non-functional properties that impact system operations – supported via a number of measurable characteristics – including but not limited to performance, security, and efficiency. As a specific example, in a resource-constrained QC platform, efficiency as a quality attribute for QCaaS ensures an efficient utilisation of the available QuBits while executing quantum algorithm as a service [S6]. In service engineering context, the concept of Quality of Service (QoS) is virtually synonymous and often used interchangeably with the concept of quality attributes. QoS refers to the overall performance and reliability of a software system as experienced by users, ensuring that it meets predefined expectations and service level agreements (SLAs) [51]. For technical distinction, we can conclude that quality attributes contribute to achieving high QoS, whereas the QoS is the actual service performance, offered by the system and/or experienced by the user. We have identified a total of six types of quality attributes including:
  • Computational efficiency refers to achieving optimal throughput and optimizing execution time etc. for scalable and high-performance computing as reported in 18 studies. In the context of computational efficiency for QCs, the term quantum computational supremacy is often used, referring to a computation threshold where a quantum computer is computationally efficient and outperforms the optimal classical supercomputers in solving a problem [2] [2]. For example, the study [S18] proposes a quantum orchestrator for load-balancing and efficient execution of quantum tasks on platforms like Amazon Braket and IBM Quantum, supporting both quantum and classical workflows.
  • Service evolution refers to the maintainability and adaptability of software services, ensuring their continuous improvement and integration. A total of seven studies have explored this aspect. To provide a more fine-grained analysis, we classify service evolution into two main types: (1) Dynamic adaptation (i.e., adaptive evolution), which enables services to adjust to changing requirements, and (2) Service integration and interoperability (i.e., perfective evolution), which focuses on enhancing compatibility and seamless interaction between services. In the context of quantum software services, service evolution is particularly critical due to the rapid advancements in quantum algorithms, hardware, and hybrid classical-quantum systems [S7]. Adaptive evolution ensures that quantum services remain efficient despite evolving quantum computing paradigms, while perfective evolution facilitates the integration of quantum services with existing classical infrastructures, promoting broader adoption and interoperability [S3].
  • Service performance measures a number of aspects such as resource efficiency, response time, energy consumption etc. to analyse and enhance the performance of quantum services. Service performance aspects can be generally classified as execution time, QuBit utilisation, and ensuring efficient quantum-classical integration, which have been identified in a total of 2 studies including [S9][S13]. For example, the study [S13] proposes a conceptual model integrating Blockchain and QC for secure, and efficient serverless edge computing where response time and resource efficiency are identified as the key criteria to measure performance of quantum services.
  • Reliability of quantum software ensures consistent execution, fault tolerance, and error mitigation, etc. to enhance dependable quantum-classical interoperability and robust quantum computing applications - identified in a total of 2 studies including [S20][S33]. As a specific example, to ensure reliability in terms of error-free execution, the study [S33] presents QuantoTrace, a cloud-based Error Correction as a Service (ECaaS) that simplifies quantum error correction for NISQ devices. The QuantoTrace solution aims to enhance quantum service reliability and accessibility while supporting future research in QC strategies.
  • Security in the context of quantum service-orientation refers to services that enable secure management and execution of systems and software executed on QC platforms - often referred to as security as a service [S14][S23]. Data security solutions offered via quantum services include robust encryption, quantum-safe cryptography, and resilience against quantum attacks, safeguarding data integrity, confidentiality, and secure communication in quantum-classical hybrid computing environments. Security as a critical quality attribute is identified in a total of 4 studies [S9][S11][S14][S23]. For example, the study [S23] proposes a smart city security architecture that can be deployed on various QC platforms, such as D-Wave and IBM Quantum Platform for real-time data classification and addressing security risks for smart city devices.
  • Usability of quantum software services focuses on intuitive interfaces, developer-friendly frameworks, and seamless quantum-classical integration, enabling efficient programming, debugging, and accessibility for researchers and practitioners in quantum computing environments, identified in a total of 4 studies [S10][S15][S30][S31]. For example, the studies [S10][S30] provide a conceptual model, focusing on a user-centric view of quantum cloud computing to enhance usability and availability of QC resources as software services.
  • Miscellaneous category includes a total of 4 studies that either lack an explicit presentation or discussion of any specific quality attributes or present one-off solutions that do not fit into any of the seven existing classification, detailed above. For example, the study [S14] explores services as sustainable software capable of leveraging QC and exploring the sustainability of QCs. The study [S5] focuses on quantum-classical microservices but does not specify any particular quality attribute.
L Quantum significant requirements represent the quantum-specific genre of the classical ASRs to ensure that functional aspects such as quantum-classic split and quality attributes such as efficient QuBit utilisation should be integral part of the conception for quantum SOAs. Functional aspects of quantum services, classified into 5 main types, mainly focus on leveraging the principles of service-orientation to deliver functional quantum services that can be invoked on QC platforms. A systematic identification and classification of key functional aspects can help to determine the features offered by quantum software services as building blocks of quantum cloud computing, a recent trend, that is gaining attention both from academic research and industrial projects to offer computational resources via pay-per-shot computing model. Quality attributes are generally classified into 7 types, mainly focusing on non-functional aspects of quantum services. In addition to supporting features as functionality, aspects such as quantum noise minimisation (service reliability), Qubit utilisation (service efficiency) are crucial for ensuring both functionality and overall quality in an attempt to support the notion of quantum computation supremacy.
Figure 5: An Illustrative Overview of the Results of RQ-2
5.2 Activity II - Service Modelling
After conception, the modelling activity relies on notations as a means to represent the design and patterns as best practices to create service design based on functional aspects and quality attributes. Modeling notations include but are not limited to graphical, descriptive, or mathematical representation of service design. Patterns complement the notations, as artifacts of reuse knowledge, to model an optimal design. As a specific example, study [S3] uses UML class and sequence diagrams as UML-based modelling notations to represent quantum search algorithms. The quantum-classic split as a layered pattern is applied to modularise the search algorithm in terms of implementing the orchestration of classic and quantum computations.
5.2.1 Modelling Notations
These are fundamental to the creation, maintenance, and evolution of models such as ontological structures and graph-based diagrams that provide a visual representation, whereas architectural description languages support a textual specification for software-intensive systems [52] [53]. Recent trends in software engineering that promote model-driven and low-code application development have resulted in transitioning developers’ focus from coding to software modeling for implementation [54]. Low-code application development process leverages the principle and practices of model-driven engineering to utilise model(s) as first-class entities in software development [54]. Moreover, with a recent emergence of large language models as software development bots, there is an increasing reliance on high-level descriptive models or graphical representation to generate source code for qunatum algorithms [55]. Investigating software models and modeling notations that help create service models is essential to support model-driven perspective to QSE, consequently facilitating quantum code developers to abstract implementation-specific complexities, via model-driven QSE, while developing quantum services [52][54]. For example, the study [S3] reports a class diagram as a structural view of the system to model the attributes and methods of entities (user, service provider, authentication etc.) of quantum computing as a service. This SMS indicates five main types of notations to model services in QCaaS detailed below based on the presentation in Figure 5.
  • Process models in service development lifecycle are used to visualise the workflows, activities, and interactions to analyze, design, and optimise services, ensuring an alignment with business goals, user needs, and requirements implemented as software services. Process models as a modeling notation are identified in a total of 3 studies [S2][S16][S17] that use a process-centric view and business process modelling notation to create service design. For example, in order to abstract out complex and implementation-specific details of quantum algorithmic implemenations, the study [S17] aims to facilitate developers to concentrate on application development rather than low-level quantum hardware details. In order to achieve this, the authors utilise the business process modeling notation, the well-known BPMN to design a process model that is used for generating quantum service source code from API specifications and quantum circuits.
  • Graph-based models have been applied in two studies using ontologies and a directed graph where the graph nodes represent a service and edges represent the interaction between the services [S7][S8]. For example, in [S7] the authors use directed graphs to represents quantum software and graph theory is used as a formal specification language of the quantum program.
  • Unified Modelling Language (UML) based models are represented via a multitude of notations, such as class and component diagrams that represent the structure [S3], while sequence and deployment diagrams [S6][S13] are used to represent the runtime or behavioural view of software services. UML diagrams and profiles represent the status-quo in software modeling and are seen as the de-facto notation in the software and service developers’ community to model classical software with growing adoption in QSE [52]. For example, the study [S10] proposes a reference architecture for quantum computing as a service. A proof-of-the-concept design and implementation of qunatum software services is provided via the Shor’s algorithm that is modeled using UML sequence diagram (modeling algorithmic behaviour) and UML deployment diagram (showing the qunatum-classic split) based deployment of the service. The role of UML-based modeling is further exemplified in the context of pattern-based modeling of quantum software services - illustrated in Figure 6.
  • Workflow model for service design represents the task to be performed (e.g., functional aspects), roles (entities or actors that interact with the tasks), and interactions to guide the structured development of software. In the reviewed studies, often the terms process modeling and workflow modeling have been used interchangeably, generally referring to modeling a flow of service execution [S16][S34]. From a technical perspective, a distinction between the two needs to be maintained. Specifically, a business process model provides a high-level view of organisational activities, focusing on goals, roles, and processes that can be automated via several services. In contrast, a workflow model details the specific sequence of tasks and data flows, often expressed as flow charts, emphasising service execution and automation. For example, the study [S34] uses workflow modeling to represent the execution and interaction of services to design, implement, and deeply qunatum algorithms.
  • Miscellaneous uses a number of different notations, often referred to as non-conventional modelling in the context of software modelling and design. These miscellaneous notations have been used in a total of 16 studies ustilising box and arrow representations of the architecture, flow charts, and mathematical notations to formally express qunatum service model. For example, the study [S19] utilises architectural overview diagram that is represented via box and arrow noation. A box represents a service in execution, whereas the arrow represents service interaction. In comparison, the study [S38] utilises mathematical notations to express QuBits and quantum circuit utilisation during service execution.
  • No exploit notation represents a scenario where no clear or explicit mention of any modelling notation for quantum services. It is vital to mention that a number of studies have used notation that are specifically developed to model qunatum software. These include qunatum circuit diagram [S10][S38] and quantum architecture description language [53].
5.2.2 Design Patterns
Patterns [^2] In software design and architecture context, the terms pattern and style are often used interchangeably to refer to best practices or frequently used solutions for recurring software modelling challenges. However, for technical clarity, a style refers to a broad structural blueprint that defines a family of systems by their components and relationships (e.g., client-server style), whereas a pattern is a reusable solution to a specific architectural problem, often addressing component interaction or deployment within a given style (e.g., a layered pattern used to implement the client-server style). represent a concentrated wisdom of software designers as reusable knowledge and best practices that can be leveraged to address recurring problems of software design and architecture. Considering a lack of professional expertise in QSE (e.g., quantum domain engineers, quantum algorithm designers, quantum software architects, etc.) patterns as artifacts of reuse can help novice developers during quantum software development to rely on existing best practices [56]. This SMS highlights that the literature on QCaaS reports five patterns, namely the API Gateway, Layered Architecture, Classic-Quantum Split, Service Façade, and Repository Pattern as shown in Figure 5. Figure 6 depicts pattern thumbnails [57] that provide a concise visual summary of the pattern, typically including its name and structural representation of pattern elements and their interactions.
Example: API Gateway for Amazon Braket enables clients to access QC resources via a Gateway - a pattern used by Amazon as a platform agnostic solution to intermediate between QC users and QC vendors via Amazon Braket [58]. For example, in Figure 6 a), the thumbnail suggests a pattern named Quantum API Gateway as a layered system comprises of three main elements including a) Quantum Service Clients that can invoke available services on b) Quantum Server via routing and recommendation by c) Quantum API Gateway, detailed in [S1]. Figure 6 a) illustrates Amazon Braket that is a managed service provided by Amazon Web Services (AWS), using the API Gateway pattern, for accessing QC resources via QCaaS. The pattern comprises three key layers. At the client node, users aim to access QC resources at the client node, having Jupyter Notebook and AWS Console, enabling them to design, simulate, and submit quantum programs via the Braket SDK. The gateway service, i.e., Amazon Braket, acts as an intermediary that abstracts hardware complexity and manages execution workflows, including device selection, queuing, and error handling. Finally, the quantum computing resources layer connects users to multiple quantum hardware providers (e.g., Vendor A, Vendor B, etc.). This model provides a platform-independent interface, facilitating seamless experimentation across various quantum technologies. By decoupling development from hardware integration, Amazon Braket supports scalable, vendor-neutral quantum computing for researchers, developers, and organisations exploring quantum applications.
Pattern documentation: One of the most common mechanisms to document and store patterns are pattern templates and pattern catalogues that provide structured management of available patterns [57]. Our intent is not to provide comprehensive details on pattern documentation and management, therefore, we rely on pattern thumbnails from Figure 6 to concisely summarise the identified patterns. We briefly introduce each pattern and elaborate on the Quantum Classic Split pattern, with an example, detailed below.
Figure 6: Thumbnail View of the Identified Patterns
  • Quantum-Classic Split is identified in a total of 3 studies is a quantum version of the Splitter pattern and enables hybrid computation that can be divided between classical and quantum computations [S10][12][s24]. As shown in Figure 6 c), the classic-quantum split is driven by the workflow, also referred to as the split logic to divide the execution of computations. Figure 6 shows that the qunatum-classic workflow modelled as a UML activity computes and delegates activities to be executed on classical and quantum computer. For example, the component and connector architectural view in Figure 6 e) shows that a ClassicalComponent and QuantumDriver deployed on a classical machine can invoke QuantumComponent on a quantum computer. A typical real-world example of the Classic-Quantum Split pattern is Shor’s algorithm for prime factorisation. The algorithm divides tasks between classical and quantum computations such that classical processing handles choosing a random integer, and performing the greatest common divisor checks, and postprocessing, whereas the quantum part efficiently determines the period using modular exponentiation and Quantum Fourier Transform (QFT). This hybrid approach leverages quantum speedup for order finding while relying on classical hardware for the remaining computations.
  • API Gateway is identified in a total of 7 studies that focuses on centralising request handling, routing, authentication, and aggregation, acting as a single entry point for microservices communication [S1][S6][S7][S15][S18][S21][S23]. For example, the study [S1] proposes the Quantum API Gateway pattern for adapting traditional API gateways for quantum services by dynamically selecting optimal quantum hardware at runtime. The study also provide a proof-of-concept implementation of the Quantum API Gateway for the Amazon Braket platform. It is vital to mention that Amazon cloud computing infrastructure, i.e., Amazon Braket uses the API Gateway pattern to manage and route requests between users and quantum computing backends [58]. It acts as an interface where developers submit quantum tasks via APIs, which are then authenticated, queued, and routed to the appropriate quantum processing unit (QPU) or simulator. This pattern abstracts backend complexities, enabling seamless integration of quantum services into classical applications.
  • Service Composition pattern combines multiple services into a single, higher-level service, enabling modularity, reusability, and streamlined functionality in distributed systems, identified in a total of 7 studies. For example, the study [S4] utilises a service composition pattern to deploy quantum services on Amazon Braket by wrapping them in classical services, demonstrating hardware variation through integer factorisation across different quantum backends.
  • Service Façade for quantum software services refers to a classical interface that abstracts and simplifies access to complex quantum functionalities. It acts as a bridge between classical clients and quantum backends, hiding the complexities of quantum computing — such as circuit design, qubit management, and backend selection — behind a well-organised and easy to implement service-oriented interfaces. Service façade is identified into a total of 2 studies. For example, the study [S7] utilises the service façade pattern, which wraps the classical (legacy software services) under the interfaces of quantum microservices that can be executed on Amazon Braket platform.
  • Miscellaneous represents a variety of different patterns such as the enterprise service bus [S16], layered architecture [S2][S23], and repository pattern [S7], identified in a total of 4 studies. For example, the study [S2] uses the layered architecture pattern to modularise a qunatum SOA via the process layer, service layer, and infrastructure layer.
L Quantum service modeling supports a model-based representation (i.e., creating design or architecture models) of functionality and quality of the conceived services. Modelling phase is supported via: – Modelling notations can assist software engineers to shift their focus from complex implementation (low-level source-coding) specific details to a design perspective (high-level system view). Modeling can incrementally transform functional aspects to service models, leading to service implementation via model-driven engineering or low-code development. The literature suggests that the modeling phase needs to leverage existing or innovative notations, such as UML diagrams and quantum circuit modells to design quantum software services. – Patterns complementăsoftware models by providing reuse knowledge and best practices of classical and quantum SE during service modelling. The review has identified several classical as well as quantum-specific patterns such as Quantum API Gateway and Qunatum-Classic Split that can facilitate architects to design quantum software services.
5.3 Service Assembly
It refers to assembling the modeled services, i.e., to implement the designed services for their deployment and execution. Service assembly phase has two main activities namely service usecase and implementation framework. Service usecase identifies the practical scenario addressed or implemented by the service, whereas the implementation framework involves tools, technologies, and programming languages etc. that enable source-coding and execution of the service. As an example, the study [S18][S19] identifies quantum simulation as a usecase that is implemented using Flask, a python-based framework for microservice-based implementation of the simulation. This section discusses the identified usecases and implementation frameworks as part of this phase.
5.3.1 Service Usecase
The literature review has identified a total of 6 use cases that include varying scenarios ranging from quantum simulation to quantum data search and security. Service usecase represents a scenario that bridges the gap between service design (modeling) and its implementation (assembly), as a practical instance of conceived services. We introduce each of the usecase below:
  • Quantum simulations is identified in a total of 4/41 studies, i.e., approximately 10% of reviewed literature [S13][S19][S20][S35]. For example, the study [S13] proposes a framework for simulation of distributed quantum computing, i.e., quantum processors that are distributed across various nodes. The study enables an experiential development, deployment, and execution of quantum software services that simulate tasks on quantum processors to evaluate the functionality and performance of distributed quantum computing.
  • Quantum search is identified in a total of 2/41 studies, i.e., approximately 5% of reviewed literature [S5][S17]. For example, the study [S] develops an experimental solution to execute quantum microservives with a usecase on searching for the best path in the context of travelling salesman problem. Quantum microservice implementation is integrated into a hybrid microservices architecture, tested on Amazon Braket across three quantum hardware vendors.
  • Algorithmic execution is identified in a total of 2/41 studies, i.e., approximately 5% of reviewed literature [S6][S24]. Execution of classical or quantum specific algorithms, such as Shor’s algorithm on quantum machines is a complex problem. For example , the study [S6] proposes a solution named quantum algorithm as a service enabling developers to abstract quantum components using software as a service and function as a service, supporting algorithm deployment across multiple quantum cloud providers and hardware platforms.
  • Cyber Security is identified in a total of 2/41 studies, i.e., approximately 5% of reviewed literature [S23][S40]. For example, the study [S23] aims to develop services to provide secure access to QC resources, ensuring that users can reliably authenticate, authorize, and interact with quantum resources while maintaining data privacy and system integrity.
  • Smart Systems is identified in a total of 1/41 studies, i.e., approximately 2.5% of reviewed literature [S11]. The study [S11] develops an experimental system to investigate quantum service-orientation in the context of smart city systems.
  • Process Automation and Optimisation is identified in a total of 4/41 studies, i.e., approximately 10% of reviewed literature [S2][S5][S7][S15]. The study [S7] proposes a hybrid quantum–classical service approach using Amazon Braket and demonstrates its applicability through a software optimization case study that tackles problems beyond the capabilities of classical algorithms.
  • No explicit usecase is identified in a total of 16/41 studies, i.e., approximately 39% of reviewed literature. For example [S14][S15] develop experimental systems but do not mention a specific usecase for implementation.
The usecase may impact the selection of programming languages and tools for service implementation. For example, the study [S3] uses Q# as the programming language that can be developed and compiled in Microsoft .Net framework for executing quantum algorithms.
5.3.2 Service Implementation
It involves programming languages that represent a system of notation or source coding scripts for implementing quantum services to manage and operationalise QC resources [25] [23]. In recent years, a number of Quantum Programming Languages (QPLs), including but not limited to Q# by Microsoft or Cirq by Google have emerged to provide specialized programming syntax and execution environment to develop, execute, and deploy quantum source code. Insights into programming languages can reveal if classical programming languages (e.g., C, Java, Python etc.) are suitable for QCaaS implementation or if there is need for more specialised QPLs (Q#, Cirq etc.). The SMS results indicate three programming languages, namely Python, Java, and Q#, as the preferred languages to implement quantum services. Based on the details of source coding, Figure 5 f) distinguishes between native code of a language and specialised libraries/application programming interfaces (APIs) being developed using a specific language. For example, the studies [S4, S5] used native Python code to implement quantum micro-servicing for experimentation. In comparison, the study [S1] used Flask as a Web framework written in Python to develop a solution for optimal delivery of quantum services on Amazon Braket. Python is the most preferred programming language both in terms of native code as well as specialised libraries of Python that include Flask and Qiskit as open-source language frameworks and Cirq which is adopted by Google. For further details on quantum programming languages, a recently conducted study [59] combines software repository mining and a practitioner survey as a mixed-method research approach to develop taxonomies of current uses and informs about emerging challenges in quantum programming as perceived by quantum developers.
  • Python as a programming language represents the predominant choice to implement quantum software services identified in a total of 19/41 studies, i.e., 46% of the total reviewed literature. Python-based service implementation can be generally classified into five sub-categories, such as native Python code [S4][S6][S35], Python Flask [S1][S5][S15], Python with Qiskit [S23][S29][S31], and Python with other languages and/or frameworks [S16][S20][S24]. For example, a quantum algorithm implemented in Python, can utilise the API Gateway pattern, for its deployment as quantum algorithm as a service on Amazon Braket platform [S6]. Python Flask is a lightweight web framework that can be used to develop web applications using Python. Flask as a web development framework for is considered as simple, flexible, and easy to get started with making it a preferred option both beginners and professionals. For example, the source code for qunatum web services is generated in Python, using the Flask Web application framework [S15].
  • Qiskit is an open-source QC framework that is developed by IBM. Qiskit is designed to support the creation, simulation, and execution of quantum algorithms [59] and identified in a total of 4 studies [S11][S21][S25][S36]. Qiskit provides tools to design quantum circuits that are used for algorithm development to manipulate quantum hardware, enabling research and development in QC. For example, the study [S36] presents a comparative analysis of qunatum software that is developed using Qiskit framework and deployed in a hybrid cloud environments.
  • Microsoft’s platform languages include several programming languages designed for developing quantum software and services that can be executed on Azure Quantum. The platform provides a comprehensive environment for quantum software development, supporting languages such as Q# and C#. Q# is specifically tailored for quantum algorithm design, while C# is used to handle classical logic and workflow orchestration. Together, they enable hybrid quantum-classical applications and seamless cloud-based execution, identified in a total of 2 studies [S3][S10]. For example, the study [S3] implements quantum function as a service for quantum application development in cloud environments using Q#.
  • Others reprents service implementation using a language and/or implementation framework not covered in any for these languages, identified in one study [S12]. The study [S12] proposes OpenAPI is a specification for defining and documenting RESTful APIs in a standard, machine-readable format. It allows developers to describe an API’s endpoints, request/response formats, authentication, and other details, typically using JSON or YAML.
  • No specific framerwork refers to the studies that do not have any explicit mention of any programming languages, as shown in a total of 15 studies. For example, the study [S2] presents a conceptual model as a model-driven development of qunatum service computing, however, no details for any implementation have been provided.
L Quantum service assembly phase focuses on assembling, i.e., implementing the modelled services, that is mainly supported by: Service usecase represents the practical context/use cases of quantum services that range from quantum search, to simulation, cyber security and proof-of-the-concept implementation of quantum algorithms. Usecass also impact the selection of programming languages and tools for implementation. Service implementation involves programming languages and/or frameworks as a system of notation for source-coding of quantum services. Classical programming languages, such as Python, represent a predominant choice over recently emerged quantum programming languages (Q#) due to more comprehensive documentation and familiarity with Python in the community of service developers.
5.4 Service Deployment
As the last phase of quantum service development lifecycle, this phase supports the selection of QC platforms on which the assembled services can be deployed for their operationalisation and execution [17]. Platform providers also referred to as quantum vendors offer computing infrastructure in terms of hardware as well as software that allows service developers to develop and/or utilise the quantum services. Deployment represents the last activity in the service life cycle that can be represented with modeling notations such as the UML deployment diagram, as shown in Figure 6. This SMS identified a total of three quantum vendors for the deployment of quantum services, namely Amazon Bracket, IBM Quantum, and Rigetti. Amazon Braket (a managed Amazon Web Services (AWS)) is the most preferred platform to design, test, and run quantum algorithms. One of the reasons for selecting Amazon Braket for service deployment is that it can allow service users/developers to design their own quantum algorithms. This can be particularly useful for novice developers unfamiliar with the technicalities of quantum systems to utilise a set of pre-built algorithms, tools, and documents to develop and manage quantum services on Amazon platform.
  • Amazon Braket is a fully managed quantum computing service by AWS that enables users to design, simulate, and run quantum algorithms using multiple quantum hardware platforms and simulators, supporting research and hybrid quantum-classical workflows in a scalable cloud environment. Amazon Braket as a deployment platform is identified in a total of 12 studies such as [S1][S4][S5][S7]. For example, the study [S1] implements a Quantum API Gateway using Python and Flask on Amazon Braket, chosen for its unified support for executing quantum code across multiple quantum processors.
  • IBM Quantum is a cloud-based platform that provides access to quantum computers and simulators. It supports Qiskit for quantum programming, enabling users to design, test, and run quantum algorithms on real quantum hardware for research, education, and development. It has been identified in a total of 10 studies, such as [S3][S10][S14][S14][S18][S19][S20]. For example, the study [S19] uses a web-based frontend for users to build and run Qiskit-based applications, with a scheduler handling job queuing, backend selection, and result retrieval across multiple execution platforms.
  • D-Wave Leap developed by D-Wave systems is a quantum annealing computer designed to solve complex optimization problems. It uses quantum annealing rather than gate-based computation, making it suitable for tasks like logistics, machine learning, and scheduling on specialized quantum hardware. It has been identified in a total of 3 studies [S11][S23][S41]. The study [S41] classical-quantum cloud architecture blueprint that spans from code-level development to production-ready quantum software applications, implemented on the D-Wave quantum platform for solving real-world optimization problems.
  • Azure Quantum is Microsoft’s cloud-based platform for quantum computing, offering access to diverse quantum hardware, simulators, and tools. It supports languages like Q# and integrates with classical systems, enabling hybrid quantum-classical application development for research, education, and enterprise solutions. It has been identified in a total of 2 studies [S10][S34]. For example, the study [S10] utilises the Miscorost Azure platform to develop a prototype for qunatum search.
  • Google Quantum AI is Google’s initiative to advance quantum computing through hardware development, algorithm research, and open-source tools like Cirq. It focuses on achieving quantum advantage and enabling scalable, fault-tolerant quantum processors for scientific and industrial applications. It has been identified in a total of 2 studies [15][16]. For example, the study [S15] provides a technical comparison of quantum computing service providers using the Travelling Salesman Problem as a case study, implemented with Cirq. Empirical tests highlight key differences and suggest approaches to reduce vendor-specific limitations, such as vendor lock-in while utilising quantum software services.
  • No Platform is represented in a total of 11 studies. For example, the study [S12] provides a conceptual model that does not include any implementation and deployment of services. Such cases of conceptual solutions lack service implementation and do not provide any details on deployment platforms.
L Quantum service deployment is enabled by deploying the implemented services on a QC server - enabling quantum service-orientation. Deployment involves selecting quantum computing (QC) platforms for executing services, marking the final phase in the service lifecycle. This study identified Amazon Braket, IBM Quantum, and Rigetti as major platforms for qunatum service deployment and delivery. Amazon Braket is the most preferred platform due to its accessible tools, support for custom algorithms, and ease of use for novice developers.
6 A Reference Architecture for QCaaS
In software engineering context, a reference architecture acts as a system-wide blueprint, defining computational elements as architectural components and representing their interactions through architectural connectors [60] [61]. The component and connector level architectural view can enable software designers to apply reuse knowledge and best practices and facilitates communication between domain professionals (a.k.a. system stakeholders) to implement the architecture [36]. Based on the results of RQ-1 detailed in Section 5 with summarised findings in Table 2, we applied the layered architecture pattern [62] to organise the service lifecycle activities into three layers namely service development, service deployment, and service split layers, illustrated in Figure 7 and detailed below. Figure 7 a) shows an abstract view to conceptualise the structural composition of the architecture, whereas Figure 7 b) uses the abstract view to present a fine-grained and concrete view of the architecture in terms of i) architectural layers, ii) phases and activities encapsulated inside each layer, iii) human roles, and iv) service artifacts. A fine-grained representation of the reference architecture, structured in individual layers and their encapsulated elements, is detailed below, as illustrated in Figure 7.
Tool support: To support architecture-based development, we have developed a prototype named ‘QADL: Quantum Architecture Description Language’ that provides a specification language, design space, and execution environment to architect quantum software services. QADL represents a proof-of-the-concept system that enables designers to specify architectural components (representing computational services), define component interconnections (enabling service communication) to support architecture-based development. Further details about QADL are provided in [53]
6.1 Layers of the Architecture
Architectural layering allows a separation of functional concerns which means architectural elements that support similar functionality can be structured in the same layer. For example, Figure 7 shows that conception, modelling, and assembly of quantum services can be unified into a layer named service development layer, service execution/hosting is presented as part of deployment layer, and the split between quantum and classic execution is presented as part of the service split layer. In addition to supporting the structuring of the system, a reference architecture provides a template that helps to design a (software) solution for a particular domain [60] - quantum computing in this case. Reference architectures also provides a common architecting vocabulary, such as functional requirements and their representation as architectural components and connectors. For example, the service development layer has one of the phases named Quantum Service Conception that encapsulates two activities namely Functional Specifications and Quality Attributes. These two activities allow a human role in the architecture such as Service Developer who can specify the required functionality and desired quality of the service. The phase produces a Quantum Significant Requirements (QSRs) as an artifact, which provides the foundation for service design. In short, architectural structuring as layers contains phases that encapsulate multiple activities, enabling the human role, i.e., the service developer to produce the service artifact for an incremental conception, modelling, assembly, and deployment of QCaaS.
6.2 Layering of Service Lifecycle Activities
The three architectural layers comprises of phases and activities of service lifecycle that are organised based on IBM service-oriented architecture (SOA) [26] to develop the reference architecture. For example, in Table 2, the phase referred to as modeling helps to focus on how to model or represent the quantum services for their implementation, while the activities named (i) modeling notation show class and component diagrams as UML-based modelling notation to enable service modelling and (ii) patterns indicate best practices to model the QC services. To architect QCaaS, we divided the Model activity from SOA life cycle into two activities, namely Conception and Modelling to distinguish between functional needs (conception) and representation (modelling) of quantum service design. The distinction allows a fine-grained representation of the reference architecture that delineates (i) the conception of functional needs and (ii) the model that architecturally represents the functional needs for their implementation. Model represents the conception as the design specification of functional needs for quantum services. For example, the initial row of Table 2 highlights that the conception, i.e., required functionality that enables the delivery of quantum software is modeled, i.e., architecturally represented using a UML deployment diagram and applying the API Gateway pattern. The reference architecture in Figure 7 do not have the Manage phase from SOA life cycle as we could not find any evidence in the literature that supports identity, compliance, and business metrics management of quantum services.
6.3 Human Roles in the Architecting QCaaS
Human roles represent human knowledge, expertise, or involvement as part of the architecting that enables development or utilisation of the quantum services [22]. There are two types of human roles in the reference architecture, referred to as service developers and service users. Service users can be individual(s) or a group of users (part of a team or an organisation) that requires quantum computation. In contrast, service developers can include a multitude of roles that include but are not limited to quantum service developers, quantum algorithm designers, and quantum domain engineers. A recent study on architecting quantum software highlights the need for quantum-specific expertise, such as quantum software architects who can map the operations of QuBits to architectural components, and quantum code managers who can simulate and analyse the flow of quantum information processing [19]. A specific human role, such as quantum domain engineer, can analyse quantum-specific attributes like mapping between QuGates and their corresponding QuBit representation. Quantum domain engineer focuses on design and fabrication activities to improve QPU performance by optimising quantum algorithms and programming languages. Quantum domain engineers, guide quantum hardware and software teams to realise quantum significant requirements that need to be implemented as quantum algorithms for execution on quantum computing platforms, as in Figure 7 b).
6.4 Service Artifacts
Service artifacts refer to specific outcomes, such as a document, design model, or source code script that enable the development and delivery of the quantum service. As in Figure 7, each phase in the architecture layer produces an artifact as an outcome of the specific phase. There are four artifacts referred to as Quantum Significant Requirements, Quantum Service Design, Quantum Service Implementation and Quantum Service Deployment. For example, the quantum service conception phase produces QSR having function and quality attributes as an artifact that supports modelling and pattern-based design of quantum service.
Figure 7: A Layered View of the Reference Architecture for QCaaS.
6.5 Example Application Scenario for the Reference Architecture
The example scenario represents a QCaaS usecase that is used to exemplify each phase and its corresponding activity in the reference architecture. The scenario does not reflect a comprehensive or full-scale development of the solution, it only elaborates a usecase of the reference architecture that provides the basis for further implementation and validation of the architecture. The example scenario as a specific instance of implementation is explained later in this section while discussing the service assembly. Implementation of the usecases is supported by QADL and detailed in our previous work [53].
  • Phase I - Service Conception: The listings below highlight the functional specification of the quantum services to generate the prime factors of a given integer. To provide this functionality, the computations need to be split between a classical machine and a quantum machine. There is also a need to assess the correctness of the implementation and efficiency of the solution for prime factorisation. The listings below as the functional aspects and quality attributes two quality attributes concerning an efficient utilisation of the QC resources in terms of the utilisation of the QuBits. The quality attributes complement the functional specification with the desired quality of the solution.
    Develop a solution that acquires an input integer N and outputs its prime factors F. - Split classical and quantum computations. - Validate correctness and efficiency.
    QuBit utilisation The solution should efficiently utilise the available qubits by minimising the number of qubits required for factoring integers of a given size.
    For example, Figure 8 indicates two architecting activities as part of service conception to specify the functional and quality aspects of prime factorisation. Functional and quality aspects are specified as quantum significant requirements that represents an architectural artifact of the conception phase. The service architect/developers can rely on the QSRs to utilise available modelling notations (e.g., UML, ADL) and apply any patterns (e.g., service orchestrator) to create a service design that acts as an artifact for assembling the microservices.
  • Phase II - Service Modeling: The QSRs, as part of the conception, act as the foundation to create a service design model that is reflective of the functional aspects. The quantum service modeling phase relies on the following two activities.
    A. Modelling notation for quantum services: As per Figure 8, we have used the UML as the modelling notation to create the service model based on UML component diagram and UML sequence diagram. To create quantum-enabled models for QSE, the quantum UML profile extends the traditional UML diagrams to support structural and behavioral modeling of quantum software [52]. Specifically, the UML component diagram in Figure 8 shows a structural view of the quantum services and their interconnection. Following the notations from UML profile [63], from a service modelling perspective each service is represented as an individual component. The component provides a computational service (e.g., generate a random number) and communicates with another component(s) using a connector. For example, the service named GetGCD interconnects with another service named Controller to generate the Greatest Common Divisor (GCD) of a randomly generated number as part of the algorithmic implementation. In contrast to a structural view of the quantum service model using a component diagram, the UML sequence diagram reflects the behavior of the services in terms of message passing among the services to enable service communication. For example, in the sequence diagram, the controller service C: Controller passes a message Generate(N) to the number generator service N: NumGenerator that generates a random number R and returns it to the controller service. A number of other UML diagrams can also be used, such as the UML use case diagram to represent the QSRs. For demonstration purposes, we have used the UML component and sequence diagrams to exemplify service modelling in terms of structural composition and behavioural representation of the services to be assembled.
    B. Pattern-based modelling: Service design and development patterns provide reuse knowledge and best practices to engineer service-oriented solutions [26]. In Figure 8, we have applied two patterns namely the orchestrator and classic-quantum split pattern. The orchestrator pattern is a classic SOA pattern that helps to orchestrate the execution of a number of services to complete a service-driven task. For example, the application of the orchestrator pattern helped to orchestrate the factorisation of a randomly generated prime number via NumGenerator and Factorise services. The classic-quantum split pattern helps to split the functionality between quantum and classic computing, also referred to as hybrid quantum computing. In the context of Figure 8, the Controller service orchestrates the functionality between classical and quantum microservices. For example, the services named NumGenerator and GetGCD can be executed on a classical machine, whereas the services QunatumModularExponentiation, QunatumInverseQFT, and Factorise are executed on the quantum machine. As part of the reference architecture, the modeling phase exploits modeling notation and applies a pattern to create a service design, i.e., a ’visual model’ that can be assembled into executable quantum microservices.
  • Phase III - Service Assembly: Service assembly refers to assembling a service, i.e., identifying a use-case and programming the service that implements the usecase as shown in Figure 8 and detailed below.
    A. Service usecase: Implementation details are elaborated based on Figure 8 that demonstrates the conception, modelling, assembly, and deployment of software services that implement the Shor’s. Shor’s algorithm is a quantum computing algorithm that is used for factoring integers in polynomial time [64]. Shor’s algorithm has a number of use cases such as prime factorisation, cryptography (e.g., breaking the RSA scheme), and finding the period of a function. In the scope of this work, we only focus on developing microservices, executed on a QC platform that implement Shor’s algorithm for prime factorisation. We provide an illustrative case in Figure 8 that exemplifies the architecture-based development of the services (from Figure 7) to implement Shor’s algorithm. Specifically, Figure 7 as a reference architecture highlights what needs to be done? by providing a blue-print to architect quantum service-orientation.
    B. Service programming: In order to implement the use-case service needs to be programmed for its execution. This means that design specifications from the modeling phase need to be translated into executable specifications as part of service assembly. Figure 8, shows the algorithmic design and its implementation as part of service assembly. The algorithm shows a partial assembly of services by exemplifying the classical services that generate a random number and compute the GCD for quantum computation of the Shor’s algorithm. The UML sequence diagram is used as a model to assemble the service via an algorithm and its implementation in a given programming language. A synoptic view of the code snippet (written in C#) is shown that highlights the source code skeleton for a controller service PrimeFactorisationController that orchestrates the classical and quantum services. Figure 8 implements the reference architecture to highlight how it is to be done? with an illustrative example of implementing quantum service-orientation.
  • Phase IV - Service Deployment as the last phase in the reference architecture, deployment consists of two main activities namely service hosting and service execution. While execution and hosting primarily depend on the QC platform on which the quantum services are deployed.
    Figure 8: Reference Architecture Based Representation of the Quantum Service Lifecycle.
    Service Hosting Platform: The hosting of the assembled service is presented as a UML deployment diagram as in Figure 8. UML deployment diagram show two deployment nodes (hosting machines). The classical compute node hosts three services NumGenerator, GetGCD, and Controller, whereas the quantum compute node hosts three services, namely QunatumModularExponentiation, QunatumInverseQFT, and Factor.
    Figure 8 b) shows the user interface of QADL, a prototype that we developed to model and assemble, i.e., represent and formally describe quantum software components that can provide computational services. In Figure 8 b), left-side, the QADL script specifies the architectural description, declaring QuBits and applying the QUGates (Hadamard, Pauli-X, and CNOT), followed by measurement. This description corresponds to the core quantum subroutine of Shor’s algorithm - superposition (via Hadamard), controlled operations for modular exponentiation (via CNOT), and state inversion (via Pauli-X) - leading to the measurement for computing the periodicity. QADL script for architectural specification of Shor’s algorithm via a quantum circuit diagram is presented in Appendix A - Listing 2.
    In Figure 8 b), towards the right, QADL represents qunatum circuit diagram to model the algorithm. The description and representation with QADL enables Shor’s algorithm to be modeled, visualised, and executed as a service abstraction that supports design at higher-level of software abstraction rather than low-level implementations. Low-level details, such as circuit creation and component execution within QADL is enabled via IBM Qiskit - an open-source QSW development framework created by IBM [65]. The Qiskit provides the tools to design, simulate, and run quantum programs on both simulators and real quantum hardware, its primary role is to provide a modular and descriptive framework for defining QSW.
7 Emerging Trends of Research on QCaaS (RQ-3)
We now answer RQ-3 that aims to report the emerging trends, as dimensions of potential future research, on engineering software services for quantum computing. To identify these trends, we specifically reviewed the details about the objectives, evaluation/demonstrations, as well as limitations and future research details in each reviewed study. This helped us to identify a number of trends that can be generally classified into five (5) main areas, detailed in this section. The identified trends presented in this section do not represent a comprehensive list, instead, it highlights some possible future research based on the data from existing literature. To maintain a consistent presentation of results, we have visualised the emerging trends identified in Figure 9 according to the service lifecycle activities [12] [26] as in Table 2. For example, Figure 9 highlights that during the conception of quantum services, an empirical identification and thematic classification of Quantum Significant Requirements (QSRs) is an area of future research. QSRs include quality aspects that complement the functional aspects to ensure the required functionality and desired quality of QCaaS as part of quantum domain engineering activity in QSE.
7.1 Process-centric and Human-assisted QCaaS Development
Process-centred engineering is manifested in service life cycle to structure a multitude of activities, such as service design, development, and delivery, into a coherent process that can be executed in an incremental fashion [22]. Although the evidence from the selected studies is accumulated and organised under SOA life cycle, however, as an individual contribution, the studies lacked a process-centric approach and highlighted the needs for processes attuned to the needs of QSE [22]. These processes, such as Quantum DevOps, Quantum microservicing or agile methods for quantum service/software etc., are seen as quantum specific genre of SE processes that align with the needs of QSE [66]. For example, in the agile method for QSE, quantum domain engineering can help accumulate the domain information of a quantum system (hardware/software) to develop a design model that can act as a blueprint to implement quantum software and services. Process-centric approaches can also support tools (for automation) and professional roles (human decision support) to engineer quantum services. We identified the need for human roles as QSE professionals to effectively undertake QCaaS development.
Process-centred quantum service-orientation can enable service developers to encapsulate qunatum source code and algorithms as modular and reusable services, designed for composition, orchestration, and execution within process-driven workflows. It can enable the assembly of end-to-end services, e.g., for optimisation, simulation, or secure cryptography, thereby shifting the focus from isolated quantum algorithms to holistic, process-aware quantum solutions.
Human roles can enrich the process - synergising knowledge of quantum physics practices of software engineering - to support activities of quantum domain engineering, software architecting, and simulation management as expertise that currently lacks in quantum service-orientation.
7.2 Empirical Identification and Classification of Quantum Significant Requirements
The concept of Quantum Significant Requirements (QSRs) can be traced to well-established role of Architecturally Significant Requirements (ASRs) in software and service engineering [48]. QSRs as part of the quantum domain engineering activity can help elicit and document the functional and non-functional aspects (also quality aspects or attributes) of quantum software. Existing research primarily focuses on functional aspects of services while overlooking the non-functional aspects that include but are not limited to QuBit utilisation, energy efficiency, quantum vendor lock-ins, QPU elasticity, quantum error mitigation, that impact the development and operationalisation or QCaaS solutions. For example, during the quantum domain engineering activity the hardware aspects (operations of QuGates) can be mapped to software aspects (components and connectors) that can help with splitting the computation tasks between classical and quantum computers using the classic-quantum split pattern.
Empirical analysis of available data available via the social coding platforms (e.g., GitHub) grounds theoretical QSRs identified via literary analysis, like Qubit utilization and error mitigation, in a practical and developer-centric context. By examining real-world projects, such as those on GitHub, research can validate or refine the identified requirements with actual developer priorities. This evidence-based approach transforms QSRs from abstract concepts into a practical taxonomy reflective of real quantum software development needs [24]. Furthermore, empirical identification and classification of the QSRs can lead to the creation of a comprehensive QSR taxonomy. By combining the findings of the literature review with data mined from open-source repositories can capture both the existing and emergent, implicit requirements from developer artifacts like issue trackers of qunatum software systems. Classifying these findings allows for the development of a taxonomical classification of QSRs and identifies critical trade-offs, such as between quantum circuit depth and quantum error mitigation [22].
Quantum significant requirements are rooted in the concept of ASRs that aims to identify, classify, and document functional and quality aspects of quantum software services. QSRs for quantum services can complement the functional aspects with quality requirements to ensure the required functionality and desired quality of service. Empirical mining of open-source social doing platforms can validate theoretical QSRs and transform them into a practical, evidence-based taxonomy reflective of real development challenges.
7.3 Model-driven Quantum Software Servicing
Model-driven Service Engineering (MDSE) enables software engineers and architects to rely on models that can abstract complex and implementation-specific details with human comprehensible visual notations to design and implement software services. Specifically, by exploiting MDSE, software engineers can apply the model transformation to transform the design models into implementation (source code) and validation (test case) models. Modeling notations, such as Service-oriented architecture Modeling Language (SoaML) and more specifically the Q-UML specification [52], provide a metamodel and a UML profile for the specification and design of services within a service-oriented architecture. MDSE can benefit novice developers to to map the flow of algorithms and modules of source code to graphical models for low-code (model-driven) development of quantum services.
Model-driven Quantum Service Development can leverage the principle of MDSE and modelling notations like SoaML and Q-UML to abstract implementation-level complexities with design-level models driven by QSRs. To exploit model-driven development, there is a need for tool support that could enable automation (e.g., model transformation tools) and human decision support (e.g., quantum software architects) to quantum software servicing using MDSE.
Figure 9: Overview of Emerging Research Trends
7.4 Empiricism in Mining Quantum Service Patterns
Pattern-based software service engineering relies on architectural design and implementation strategies and best practices that can be reused to deliver software services [26]. Existing solutions employ a number of pattern-based solutions, such as classic-quantum split and service wrapping pattern, however, there is no evidence on an empirical discovery of patterns and tactics as reusable knowledge [67]. A lack of empiricism in discovering new patterns hinders the reusability of service design and implementation knowledge during quantum service engineering. One possible dimension for pattern discovery is mining software repositories or social coding platforms (e.g., GitHub) that contain raw knowledge that can be mined as patterns. Pattern-based solutions could complement human expertise with available best practices for service design and implementation. More specifically, novice developers of quantum software services can rely on patterns that abstract complexities of quantum source coding with modularisation of pattern-based solutions.
Pattern discovery via empirically-grounded methods, i.e., mining repositories or social coding platforms of quantum software development, can help leverage reusable design rationale for quantum service engineering. Pattern languages can empower the role of quantum software architects and algorithm designers to rely on reusable knowledge and best practices as opposed to ad-hoc and once-off solutions.
7.5 Continuous Testing and Delivery of Quantum Services
With an adoption of agile software engineering in quantum software development context [66], there is a need for light and adaptive methods to ensure a continuous development and delivery of quantum software services. The literature suggested a lack of solutions on testing the software services. The continuous testing and continuous delivery can help CT/CD to test the services against QSRs more effectively and deliver them rapidly. Quantum service testing can involve simulation or regression tests against the QSRs. Quantum service testing involves various approaches, including simulation and regression tests against QSRs. Given the unique characteristics of quantum computing, such as superposition and entanglement, traditional testing methods may not suffice. Therefore, specialised testing strategies are essential to ensure the reliability and performance of quantum services.
Furthermore, integrating CT/CD practices into the quantum software development lifecycle facilitates early detection of issues, reduces feedback times, improves visibility of product quality, and accelerates time-to-market [68]. This integration is crucial for maintaining the agility and adaptability required in the rapidly evolving field of quantum computing
Continuous Testing and Continuous Delivery (CT/CD) relies on agile software engineering methods to help deliver quantum software services rapidly and reliably. CT/CD can provide strategic benefits to vendors by adding new services to their quantum platforms.
Large Language Models and Quantum Service Engineering: Although we did not find any explicit evidence, specifically published research on synergising generative artificial intelligence (including the large language models (LLMs)) and quantum service development. However, some recent research studies indicate a growing trend on exploiting LLMs to design and develop quantum software. Recent work shows that LLMs can mimic the behaviour of quantum circuits and predict how quantum states evolve, thus enabling a faster and resource efficient alternative to traditional quantum simulators [69]. Purpose-built LLMs such as the GroverGPT, have claimed to achieve near-perfect accuracy when replicating Grover’s search algorithm on systems with up to 20 qubits [70]. Beyond the general purpose simulations, LLM-based solution are improving the quality of quantum source code through tools such as LintQ-LLM, which can detect, locate, and explain programming errors in Qiskit more effectively than standard static analysis methods [71]. They are also being used to help developers keep up with rapid changes in quantum programming libraries—such as migrating between Qiskit versions—by automatically identifying and classifying common refactoring scenarios [72]. The collective impact of these advancements indicates that LLMs, already recognised for their role in software engineering and architecting [73], are likely to become central to quantum software engineering—enabling code generation and testing, as well as supporting the simulation, development, and maintenance of quantum software services.
QCaaS in the era of LLMs: Recent studies show LLMs can simulate quantum circuits, enhance code quality, and support library migration in quantum software development. While direct research on combining LLMs with quantum service computing is scarce, emerging tools like GroverGPT and LintQ-LLM indicate LLMs may soon become central to quantum software engineering.
8 Related Survey-based and Empirical Studies
We now review the most relevant existing research in terms of survey-based research on (i) quantum software engineering (Section 8.1), and (ii) quantum services computing (Section 8.2). The discussion in this section is complemented with data in Table 3 that acts as a structured catalogue for criteria-based comparative analysis of proposed vs existing research.
Interpretation of the Comparative Analysis: For an objective comparison, Table 3 presents each study using five-point self-explanatory criteria including (a) type of study (adopted from [21]), (b) focus of study, (c) core findings, (d) QSE activity supported by the study, and (e) year of publication, each exemplified below. For example, the study [19] presents a systematic review, conducted in 2021, as part of evidence-based software engineering to focus on architecting quantum software. The SLR presents the core findings about architectural life cycle and state-of-the-art on architectural modeling, patterns, and tools to architect quantum software. The SLR focuses on design and architecting activities of QSE life cycle and indicates the needs for future research on quantum-specific professional expertise (human roles in QSE) to develop quantum software.
Table 3: A Review of the Relevant Secondary Studies
Quantum Software Engineering
Study
Reference
Type of
Study
Focus of
Study
Core
Findings
QSE
Activity
Publication
Year
[35]
Literature
Review
Quantum
Software
Engineering
QSE life cycle, quantum software
engineering processes, methods, and tools.
Software
Life cycle
2020
[19]
Systematic
Review
Quantum
Software
Architecture
Quantum software architecting activities,
modelling notations, patterns, and tools
for architectural development
Software Design
and
Architecture
2021
[25]
Repository Mining
and
Practitioner Survey
Quantum
Programming
Languages
Mining repositories and interviewing
practitioner to investigate quantum
programming languages
Software
Implementation
2022
[74]
Repository
Mining
Quantum
Issues
Mining repositories to identify technical
debts in open-source quantum software
Software
Implementation
and Testing
2022
[75]
Vision
Paper
Quantum
Software
Architecture
QSE life cycle and comparing quantum
computers with their classical
counterparts and vision for future research
Software
Life cycle
2022
[76]
Systematic Mapping
Study
Quantum
Software
Engineering
Mapping QSE research, identifying
key topics, study types, results,
popular quantum tools, and the
evolution of community interest
Facilitate knowledge transfer
(i.e., development and evolution,
the tools and frameworks
being studied, etc.)
2024
[18]
Literature
Review
Security challenges
associated
with QSE
Addressing security in quantum
programming languages
, hardware limitations, and
their impact on software
development, highlighting vulnerabilities
and advocating for early integration
of security measures to prevent exploitation.
Software Design and
Testing
2023
[77]
Literature
Review
Quantum Software
Engineering
Investigating the evolving
landscape of QSE, synthesising
insights to highlight key
challenges and the need
for robust frameworks and
methodologies to advance the of QSE.
Software
Lifecycle
2024
Quantum Services Computing
[16]
Literature
Survey
Quantum
Cloud
Computing
Programming quantum computers and
investigating hybrid software consisting
of classical parts and quantum parts.
Service Design
and
Architecture
2020
[78]
Literature
Review
Quantum
Cloud
Computing
Reviewing advances, challenges,
and future directions in
quantum cloud computing,
covering models, platforms,
technologies, and key
issues (e.g., resource
management, security, and privacy).
Identifying open problems
and highlights opportunities
for future research.
Service Lifecycle
2024
[79]
Literature
Review
Distributed Quantum
Computing
Discussing the challenges and open
issues in distributed quantum computing
from a computing and communications
perspective.
Distributed Quantum
Computing (Algorithms, Compiling,
Networking and
Simulation)
2024
[80]
Literature
Survey
Quantum
Service
Computing
Potential and limitation of integrating
quantum computing with cloud computing
Service
Lifecycle
2015
8.1 Quantum Software Engineering
Quantum Software Engineering (QSE) has emerged as the most recent genre of software engineering (SE) that allows practitioners to adopt a process-centric and systematic approach to develop software-intensive systems and applications for QC [35, 75]. We have found a total of 5 survey-based studies, summarised in Table 3 to compare the scope and contributions of our proposed research. Some recently published survey-based studies on QSE highlight that existing SE principles and practices can be tailored to develop quantum software, however, issues specific to QC, such as operationalising QuBits/QuGates, quantum domain engineering, and quantum-classic software split, require new engineering methods to tackle these challenges [75]. To derive new methods and processes, the QSE research community is striving to organize a body of knowledge and academic collaborations that can be stimulated via dedicated publication forums to disseminate quantum software specific research including, but not limited to IEEE QSW  [^3] IEEE International Conference on Quantum Software. https://services.conferences.computer.org/2025/qsw/qsw-call-for-papers/ and ACM TOSEM’s Open Continuous Special Section on Quantum SE [^4] ACM Transactions on Software Engineering and Methodology. https://dl.acm.org/journal/tosem/quantumse . Academic discussions at the publication fora and published results [1][4] have highlighted that in addition to the engineering principle and practices, human knowledge and expertise require fundamental knowledge of quantum mechanics to effectively design algorithmic solutions for QSE projects [25]. The current generation of software architects and developers who lack the foundational knowledge of quantum mechanics may be hindered or find themselves under-prepared for quantum software development [19]. Software architecture is being viewed as a solution that can abstract complexities of implementation by representing software to be developed as architectural components and connectors. A recently conducted systematic review of architectural solutions puts forward five architecting activities to guide software designers to engineer quantum software solutions in an incremental manner [19]. Architectural solution(s) can also empower quantum software engineers to exploit architectural knowledge in terms of patterns (i.e., reuse knowledge), modelling notations (i.e., software representation), and tools (i.e., automation) to address the challenges of QSE. In addition to architecting, there is a need for empirically grounded guidelines and solutions to address the aspects of quantum software programming, model-driven development, and quality assurance in QSE [74].
Despite the recent advancements in Quantum Software Engineering (QSE) discipline, a multitude of technical and human skills related challenges persist that require innovative solutions and empirical research to address them. The study in [76] presents the results of a systematic mapping study of the current state of QSE research and identifies trends such as the most investigated topics, types and number of studies, main reported results, and the most studied quantum computing tools and frameworks. The study concludes by exploring the research community’s interest in QSE, how the discipline is evolving, and implications of research results are as a centralised source of information to facilitate knowledge transfer for researchers and practitioners. The authors of [81] have conducted a two-phase empirical study to investigate the impcat of qunatum programming language on qunatum software development. Specifically, at the first stage, the authors mined data from GitHub repositories to investigate the quantum programming frameworks used by developers for qunatum software development. Secondly, the authors conducted a survey targeting the contributors of these repositories to capture their perspectives on the current adoption and challenges of quantum programming. The study [20] conducted in-depth semi-structured interviews to investigate practationers’ view on the issues and challenges of quantum software engineering. The insights from interviews provide developers’ perspectives such as the motivations, challenges, and outlooks within the rapidly evolving field of QSE. Despite promising advancements in QSE, significant challenges persist and demand attention. De Stefano et al. in [76] analysed widely-used quantum tools, offering valuable insights into their capabilities and limitations. Arias et al. in [18] explored security issues in quantum programming languages and hardware, emphasising their impact on software development. The study highlighted critical vulnerabilities and advocated for the early integration of security measures to mitigate potential exploitation. Murillo et al. [77] further emphasised the need for robust frameworks and comprehensive security protocols. As quantum computing continues to advance, integrating QSE into practical applications becomes increasingly critical, requiring a thorough understanding of both its potential and associated risks.
8.2 Quantum Service Computing
Quantum service-orientation is a recent trend initially pushed by QC vendors to allow developers who can compose and invoke software services on remotely hosted quantum systems and infrastructures [16]. We identified a total of 2 survey-based studies on quantum service computing highlighted in Table 3. Quantum service computing initiatives, such as Amazon Braket and Q experience, have paved the way for academic research to propose solutions for quantum cloud, quantum services computing, and quantum as a service. Recent research reviews inform about the current progress and emerging challenges for quantum services computing that include but are not limited to hardware availability, quantum noise, and quantum-classic split [80]. Software researchers are striving to synergise existing methods of microservicing and quantum software development, such as Quantum DevOps to develop quantum microservices [66].
Researchers on quantum cloud computing and distributed quantum computing highlighted both significant advancements and ongoing challenges in leveraging quantum technologies for practical use. Both fields emphasised the importance of integrating quantum computing with existing infrastructures to expand computational capabilities and overcome current limitations. The study in [16] provided a comprehensive review of quantum cloud computing, examining models, platforms, technologies, and critical issues such as resource management, security, and privacy. Their study identified open problems and outlined opportunities for future exploration. Similarly, the study [80] explored the challenges and unresolved issues in distributed quantum computing, offering insights from a computing and communications perspective. In addition, researchers on quantum cloud computing (QCC) and distributed quantum computing highlighted both significant advancements and ongoing challenges in leveraging quantum technologies for practical use. Both fields emphasised the importance of integrating quantum computing with existing infrastructures to expand computational capabilities and overcome current limitations. Nguyen et al. [78] provided a comprehensive review of quantum cloud computing, examining models, platforms, technologies, and critical issues such as resource management, security, and privacy. Their study identified open problems and outlined opportunities for future exploration. Similarly, Caleffi et al. []explored the challenges and unresolved issues in distributed quantum computing, offering insights from a computing and communications perspective.
9 Implications and Validity Threats
This section discusses the potential implications of study results on researchăand practice of quantumăcomputing as a service in Section 9.1. We also present some threats to the validity of the findings from this study in Section 9.2.
Figure 10: Synergising Quantum and Service Computing
9.1 Implications of Study Results
This research focuses on investigating the potential synergy between quantum computing and quantum service computing in a rapidly emerging domain of quantum service computing, as illustrated in Figure 10. The study results synthesise software engineering principles applicable to QCaaS, defines service lifecycle phases in a reference architecture as a blueprint for quantum service-oriented development, and highlights key trends in terms of processes, professional expertise and tool support, etc. to guide future research and advance next-generation QCaaS solutions. We highlight the implications of the study in terms of potential impacts of the results as valuable insights for both research and practice on quantum software engineering in general and quantum service computing in particular.
9.1.1 Implications for Research
The results of this study are presented as structured catalogues (Table 2) and illustrative visualisations (Figure 2 - Figure 7) that provide a synthesized knowledge on research state-of-the-art and vision for future needed research. Core findings of the study organised as service lifecycle activities in Section 5 and their integration into a reference architecture in Section 6 can inform researchers about leveraging existing SOA principles, practices, and patterns in the context of emerging design challenges of QCaaS. Specifically, the reference architecture derived from existing literature provides a blue-print that harnessesăpatterns and tactics from classical software engineering (layered architecture, service facade) that can be integrated with quantum-specificăpatterns such as the quantum API gateway and quantum-classic split. An illustrative discussion of emerging trends can provide a roadmap as in Figure 9 for future research involving empiricism such as surveys, interviews, or repository analyses research to develop evidence-based taxonomies to architect and implement QCaaS. The study results also indicate that the ‘Manage phase of the SOA lifecycle [26] is currently underexplored, offering potential for future research on quantum-specific service monitoring, governance, and quality management. Researchers could explore how LLMs might automate tasks such as generating quantum service models from natural language, completing quantum code, migrating code between frameworks, or simulating quantum circuits - speeding up the development via human-bot collaborative software engineering for QCaaS.
9.1.2 Implications for Practice
The results of this study provide evidence-based guidelines for practitioners, including quantum software architects and developers about the QSRs that guide quantum software and service development process (Section 8). The pattern catalogue and thumbnails in Figure 6 represent a concentrated wisdom of quantum service that enables software designers and architects to rely on reuse knowledge and best practices to engineer QCaaS. The study indicates that QSE requires dedicate industrial roles such as qunatum domain engineer, quantum simulation manager, quantum algorithm desinger in an industrial scale development of unatum software as in Figure 9. The predominance of Python-based programming frameworks and qunatum service execution environments such as Qiskit and Amazon Braket reflects the gradual maturity quantum software/service ecosystem, reduces adoption risk, and enables practitioners to leverage established guidelines and concrete frameworks for service assembly and deployment. The study can help practitioners with the structured knowledge, patterns, and architectural guidance needed to build reliable and efficient quantum services, helping to enable and enhance software engineering for QCaaS.
9.2 Threats to Validity
We now discuss several threats that may constrain or limit the validity of a study’s findings [82]. These threats are commonly grouped into three main categories, as outlined below.
9.2.1 Threats to Internal Validity
Internal validity addresses whether the results of the study (e.g., service lifecycle phases, and derived reference architecture) are influenced only by the variables under study (e.g., SMS process, ABD method). Potential threats to internal validity may arise from factors including, but are not limited to literature selection bias, the subjectivity of quality assessment criteria, and inconsistencies in data extraction that have been presented in Section 3. We tried to minimise these threats by following a structured process (Figure 3), specifically customising the search string from Listing 1 as per individual EDS by following the guidelines in [41]. Moreover, to ensure a multistep quality assessment of each individual study, we derived a qulaity assessment checklist, as presented in Table 1. Finally, to minimise the threat of omitting key information from the literature, we created a structured template as the data extraction form Appendix A - Table 4 representing 11 data points to extract and record data from each primary study. The data extraction template provided a foundation to synthesise the data and document the study results.
9.2.2 Threats to External Validity
External validity concerns the extent to which the study’s findings can be generalised beyond the immediate scope of this research, including to different quantum platforms, service models, or development environments. These threats can include the dominance of academic research over industrial perspectives/solutions, temporal restrictions of the reviewed literature, and the narrow evidence base, i.e., 41 studies as the underlying evidence for the proposed reference architecture. As the findings are derived primarily from peer-reviewed academic sources (Appendix A - Table LABEL:tab:study-list), they lack practitioners’ perspective and industry-based insights for quantum service architecting and development. One possible way to minimise this threat is by using a mixed-method approach [83] that uses findings from literature to derive an architecture and develops a proof-of-the-concept validation as in Figure 8. The limitation can be further mitigated with future work that relies on survey research with quantum service developers and software engineers to validate and complement the findings of the study. Furthermore, extending the evidence base through approaches such as repository mining or practitioner-oriented evaluation of the reference architecture can provide stronger empirical grounding and enhance the generalisability of the proposed architecture.
External validity could also be affected by study selection, i.e., publication bias. We adhered to the guidelines of systematic reviews [39] and only included the peer-reviewed published literature for data extrcation. By focusing primarily on peer-reviewed literature and excluding most gray or multivocal literature [33] such as preprints, this study may underrepresent findings such as the service lifecycle phases and activities in each phase. While the decision to include only peer-reviewed published research enhances methodological rigour (per the guidelines of SLRs and/or SMS), it potentially limits the representativeness of available evidence as gray literature to limit the findings of this research.
9.2.3 Threats to Construct Validity
Construct validity addresses the extent to which a study accurately measures what it is intended to measure, ensuring that the research questions, methodological details, etc. are clearly defined and appropriately operationalised. In the context of this study, construct validity relates to whether the selected research questions, study selection criteria, and data extraction items (in Section 3) adequately capture the essential aspects of Quantum Computing as a Service (QCaaS) and its software engineering underpinnings. A key threat arises from the operationalisation of core constructs such as quantum service lifecycle, reference architecture, and quantum significant requirements (QSRs). Although these constructs have been defined based on established software engineering and quantum computing literature [35] [19], the rapid evolution of terminology within the quantum software engineering domain introduces a risk of conceptual drift, potentially limiting the generalisability of our findings to future research contexts.
Another potential threat is the subjectivity in the classification and coding of primary studies. Despite employing a structured data extraction form, interpreting constructs such as ‘pattern’, ‘modelling notation’, or ‘service quality attribute’ in Table 2 can be influenced by the authors’ judgment. To mitigate this threat, we adopted a multi-step verification process wherein the first author performed data extraction, and subsequent verification was independently conducted by co-authors.
10 Conclusions
Quantum service orientation allows QC vendors to offer end-users’ access to computational resources - enabling shots at QPUs - by following the utility computing model. Recently, research and development on QSE have started to synergise the principles of service-orientation and practices of QC (algorithms, simulations, QuGates, etc.) to enable the adoption of QCaaS by individuals and enterprises to attain quantum supremacy in modern day computing. To this end, we used the systematic mapping study approach to primarily investigate (i) existing solutions for quantum service development, such as QSRs, modelling notations, patterns, and programming frameworks etc. to architect QCaaS, and (ii) emerging trends that highlight the needs for ongoing and future research on QCaaS. The results of this SMS as structured information (Table 2) and visual illustrations (Figure 2, 5 - Figure 7) highlight the strengths, limitations, and future research from QSE perspective.
The primary contributions of this research focus on synthesising the existing research-based evidence on software engineering for QCaaS, proposes a reference architecture to guide quantum service lifecycle, and highlights emerging research trends on quantum service orientation. The findings and implications of the study can guide researchers with an evidence-based map of QCaaS and provide practitioners with patterns as reuse knowledge and architecture as a blue-print to implement solutions.
Needs for future work: The primary implication of this study is to establish an evidence-based body of knowledge for service development that leverages design notations, patterns, and architectural approaches highlighting the needs for human-centric and process-driven QCaaS development. The results of this SMS establish the foundations for future work in three directions that include (a) conducting practitioners’ survey, (b) implementing the reference architecture, and (c) mining social coding platforms for empiricism in QCaaS computing. The literary foundations can help us to design an empirical study that aims to engage service developers and engineers in a workshop (activity and survey) to seek their feedback and synthesise the results as practitioners’ perspectives to complement the evidence from academic research. We also plan to mine social coding platforms (e.g., GitHub) to empirically discover knowledge and understand the practices adopted by developers’ communities in open-source QCaaS.
[…]
Appendix A Appendix
Listing 2: QADL Script for Specification of Shor’s Algorithm
@startqadl
Circuit ShorsAlgorithm {
qubit q0, q1, q2
gate Hadamard q0, q1, q2
gate X q0 // Invert q0
gate X q2 // Invert q2
gate CNOT q0 q1 // Controlled q0, q1
gate CNOT q1 q2 // Controlled q1, q2
gate X q0 // Undo inversion of q0
gate X q2 // Undo inversion of q2
gate Hadamard q0, q1, q2
gate X q0, q1, q2
gate CNOT q0 q1
gate CNOT q1 q2
gate X q0, q1, q2
gate Hadamard q0, q1, q2
measure q0 -> c0
measure q1 -> c1
measure q2 -> c2
@endqadl
Table 4: Data Extraction Template
Generic Data
D1 Study ID The unique identifier for each study for its reference in the SMS
D2 Study Title The title of the study
D3 Study Authors The authors as contributors of each study
Demographic Data of Published Studies (RQ-1: Research Demography)
D3 Year of publication The specific year in which a study is published to identify the overall temporal distribution of publications
D4 Type of publication The type of research publications to identify publication foras such as conference proceedings or journal articles etc.
D5 Type of Research Facet The type of research contribution to identify the strength of published evidence such as solution proposal, vision paper etc.
D6 Application Domain The type of system(s) or scenario(s) such as cyber security, smart city etc. that is supported by the proposed research
Data for Quantum Service Development Lifecycle (RQ-2: Quantum Service Lifecycle)
D7 Conception The data for conceptualising the quantum service
a Functionality Functional requirements as the conceived functionality offered by the service
b Quality Quality attributes as the conceived quality of the service
D8 Modeling The data to model, i.e., specify or represent the conceived service
a Notation Means to represent the design of the conceived service
b Pattern Reusable knowledge or recurring solution(s) as applied best practices to design the service
D9 Assembly The data to assemble, i.e., program or implement the modeled service
a Usecases The scenario that is implemented or automated by the assembled service
b Programming Programming language or technologies used to assemble/implement the service
D10 Deployment The data to deploy the assembled service
a Delivery The quantum computing platform that is used to deploy and enable the execution of the deployed service
Data for Emerging Trends and Potential Future Research
D11 Emerging Trends Identify the emerging trends such as open challenges and/or needed solutions as potential dimensions of future research
Table 5: List of Primary Studies for the SMS
Study ID Study Title Publication Venue Year
S1 Quantum software as a service through a quantum API gateway IEEE Internet Computing 2021
S2 QSOC: Quantum service oriented computing Symp. and Summer School on Service-Oriented Computing (SummerSOC) 2021
S3 QFaaS: A Serverless Function-as-a-Service Framework for Quantum Computing Future Generation Computer Systems 2024
S4 Quantum service-oriented computing: current landscape and challenges Software Quality Journal 2022
S5 Trials and tribulations of developing hybrid quantum-classical microservices systems Quantum Software Engineering and Technology Workshop (QSWE) 2021
S6 Towards Quantum algorithms-as-a-service Int. Workshop on Quantum Programming for Software Engineering (Q-SE) 2022
S7 Quantum Service-Oriented Architectures: From Hybrid Classical Approaches to Future Stand-Alone Solutions Quantum Software Engineering 2022
S8 Relevance of near-term quantum computing in the cloud: A humanities perspective Int. Conf. on Cloud Computing and Services Science (CLOSER) 2020
S9 Hybrid classical-quantum software services systems: Exploration of the rough edges Int. Conf. on the Quality of Information and Communications Technology (QUATIC) 2021
S10 A reference architecture for quantum computing as a service Journal of King Saud University – Computer and Information Sciences 2024
S11 Securing Smart Cities: Unraveling Quantum as a Service Int. Workshop on Quantum Programming for Software Engineering (Q-SE) 2023
S12 QCSHQD: Quantum computing as a service for Hybrid classical-quantum software development: A Vision ACM Int. Workshop on Quantum Software Engineering (QSE-NE) 2024
S13 A Design Framework for the Simulation of Distributed Quantum Computing Workshop on High Performance and Quantum Computing Integration (HPQCI) 2024
S14 Defining Quantum Advantage for Building a Sustainable MVP to Deliver Quantum Computing Services Open Journal of Applied Sciences 2024
S15 Development and Deployment of Quantum Services Quantum Software: Aspects of Theory and System Design 2024
S16 Integration of Classical and Quantum Services Using an Enterprise Service Bus Int. Conf. on Product-Focused Software Process Improvement (PROFES) 2023
S17 Enabling continuous deployment techniques for quantum services Journal of Software: Practice and Experience 2024
S18 Orchestration for quantum services: The power of load balancing across multiple service providers Science of Computer Programming 2024
S19 Qubernetes: Towards a unified cloud-native execution platform for hybrid classic-quantum computing Information and Software Technology 2024
S20 Quantum Service-Oriented Computing: A Practical Introduction to Quantum Web Services and Quantum Workflows Int. Conf. on Web Engineering (ICWE) 2024
S21 Quantum services generation and deployment process: a quality-oriented approach Int. Conf. on the Quality of Information and Communications Technology (QUATIC) 2023
S22 Quantum computing in the cloud: Analyzing job and machine characteristics IEEE Int. Symposium on Workload Characterization (IISWC) 2021
S23 Quantum as a Service Architecture for Security in a Smart City Int. Conf. on the Quality of Information and Communications Technology (QUATIC) 2023
S24 Quantum web services: Development and deployment Int. Conf. on Web Engineering (ICWE) 2023
S25 Quantum microservices: transforming software architecture with quantum computing Int. Conf. on Advanced Information Networking and Applications (AINA) 2024
S26 Evolution of Service-Oriented Computing: Integrating Quantum Techniques for Integer Factorization IEEE Int. Conf. for Convergence in Technology (I2CT) 2024
S27 Quantum and blockchain based Serverless edge computing: A vision, model, new trends and future directions Internet Technology Letters 2024
S28 Enabling continuous deployment techniques for quantum services Journal of Software: Practice and Experience 2024
S29 Quokka: a service ecosystem for workflow-based execution of variational quantum algorithms Int. Conf. on Service-Oriented Computing (ICSOC) 2022
S30 Quantum Cloud Computing from a User Perspective Int. Conf. on Innovations for Community Services (I4CS) 2023
S31 Towards a Quantum-Science Gateway: A Hybrid Reference Architecture Facilitating Quantum Computing Capabilities for Cloud Utilization IEEE Access 2023
S32 Hybrid Quantum Machine learning using Quantum Integrated Cloud Architecture (QICA) Int. Conf. on Computing, Networking and Communications (ICNC) 2023
S33 QuantoTrace: Quantum Error Correction as a Service for Robust Quantum Computing Int. Conf. on Electrical Engineering and Information & Communication Technology (ICEEICT) 2024
S34 A Workflow for the Continuous Deployment of Quantum Services IEEE Int. Conf. on Software Services Engineering (SSE) 2023
S35 A quantum computing simulator scheme using MPI technology on cloud platform IEEE Int. Conf. on Electrical Engineering, Big Data and Algorithms (EEBDA) 2022
S36 Comparing the Orchestration of Quantum Applications on Hybrid Clouds IEEE/ACM Int. Symp. on Cluster, Cloud and Internet Computing Workshops (CCGridW) 2023
S37 Dynamic Quantum Network: from Quantum Data Centre to Quantum Cloud Computing Optical Fiber Communication Conference (OFC) 2022
S38 Stochastic Qubit Resource Allocation for Quantum Cloud Computing IEEE/IFIP Network Operations and Management Symp. (NOMS) 2023
S39 Architectural Vision for Quantum Computing in the Edge-Cloud Continuum IEEE Int. Conf. on Quantum Software (QSW) 2023
S40 A Quantum Safe Approach for Security Challenges at the Edge of Cloud in 5G and Beyond IEEE Cloud Summit 2024
S41 Quantum Software Architecture Blueprints for the Cloud: Overview and Application to Peer-2-Peer Energy Trading IEEE Conf. on Technologies for Sustainability (SusTech) 2023